### v1

# Self-Guided Open-Vocabulary Semantic Segmentation

Abstract
Vision-Language Models (VLMs) have emerged as promising tools for open-ended image understanding tasks, including open vocabulary segmentation. Yet, direct application of such VLMs to segmentation is non-trivial, since VLMs are trained with image-text pairs and naturally lack pixel-level granularity. Recent works have made advancements in bridging this gap, often by leveraging the shared image-text space in which the image and a provided text prompt are represented. In this paper, we challenge the capabilities of VLMs further and tackle open-vocabulary segmentation without the need for any textual input. To this end, we propose a novel Self-Guided Semantic Segmentation (Self-Seg) framework. Self-Seg is capable of automatically detecting relevant class names from clustered BLIP embeddings and using these for accurate semantic segmentation. In addition, we propose an LLM-based Open-Vocabulary Evaluator (LOVE) to effectively assess predicted open-vocabulary class names. We achieve state-of-the-art results on Pascal VOC, ADE20K and CityScapes for open-vocabulary segmentation without given class names, as well as competitive performance with methods where class names are given. All code and data will be released.
1 Introduction
Semantic segmentation is a computer vision task that involves grouping pixels into coherent regions based on their semantic categories. While notable advancements have been made, existing semantic segmentation models primarily rely on labeled datasets with predefined categories, limiting their ability to handle unknown classes. In contrast, humans possess an open-vocabulary understanding of scenes, recognizing thousands of distinct categories. Recent studies have focused on the development of segmentation models to meet similar capabilities [21, 18], with models that leverage Vision-Language Models (VLMs) as an emerging category [14, 10, 7, 16, 5, 36, 3, 34, 35]. Such VLMs learn rich multi-modal features from billions of image-text pairs.
Figure 1: Self-Guided Open-Vocabulary Semantic Segmentation. Our method is capable of determining and segmenting relevant open vocabulary class names automatically.
Given the scale, training VLMs comes at a extremely high computational cost. Because of this, it is common in literature to build upon pre-trained VLMs such as CLIP [22] or BLIP [15]. However, applying these VLMs on per-pixel tasks is non-trivial, as they are trained on full images and lack the ability to directly reason over local regions in the image. Previous attempts have employed a two-stage approach where masked regions are fed to the VLMs to produce CLIP embeddings, which are subsequently correlated with text encodings of text prompts to perform classification [33, 16, 7]. Other works directly align per-pixel embeddings with text embeddings retrieved from CLIP [14] based on text prompts provided by the user at test time. While training these models, the text prompts usually correspond to the ground truth class names of the dataset in question. In this paper, we take the capabilities of VLMs one step further and tackle a more challenging task in which the relevant class names are automatically generated. We enable self-guided open-vocabulary semantic segmentation, in which the goal is to produce accurate object segmentation and their pixel-level classifications for any class without the need for textual input from the user or predefined class names.
To tackle this task, we introduce a Self-Guided Semantic Segmentation (Self-SEG) framework. Our framework can adapt to any pre-trained segmentation backbone limited by the need for textual input describing the class names to be potentially segmented. In order to automate this, we propose BLIP-Cluster-Caption (BCC) which consists of four simple steps: 1) obtain local BLIP embeddings at different image scales; 2) cluster local BLIP embeddings to identify and separate meaningful semantic areas in the image; 3) caption each clustered embedding and extract nouns from the caption; and 4) pass the resulting nouns to a pre-trained segmentation backbone as class names, thereby essentially creating a self-guided network without any additional supervised training.
Experimental evaluations on Pascal VOC [9], ADE20K [37] and CityScapes [6] showcase the effectiveness of our framework. Self-SEG achieves state-of-the-art results on self-guided open-vocabulary semantic segmentation and competitive results with methods where class names are given. It is capable of predicting crisp segmentation masks without having any information a priori about which objects to segment (see Fig. 1). For effective evaluation on public datasets, we furthermore propose an LLM-based Open-Vocabulary Evaluator (LOVE), tasked with mapping open-vocabulary predictions to class names in the dataset using a Large Language Model (LLM). All code will be publicly released to ensure reproducibility and to facilitate further research.
In summary, our contributions are as follows: 1) we introduce Self-SEG, a novel framework to automatically determine and segment relevant classes in an image; 2) we propose BLIP-Cluster-Caption, a novel image captioning method capable of generating local and semantically meaningful descriptions from the image obtained from clustered BLIP embeddings; and 3) we propose LOVE, a novel evaluation method for open-vocabulary semantic segmentation which utilizes an LLM.
Through qualitative and quantitative analyses, we demonstrate the effectiveness of our Self-Guided Semantic Segmentation framework on multiple public datasets.
Figure 2: Different Semantic Segmentation Tasks in Comparison. Left: In traditional semantic segmentation, an image is segmented into fixed, predefined dataset classes used during annotation. Middle: In open-vocabulary segmentation, the user specifies which object categories (from the open vocabulary) need to be segmented. Right: In self-guided segmentation, relevant object categories are generated from the image automatically, enabling a true open-vocabulary understanding of scenes without a human in the loop.
2 Related Work
Open-Vocabulary Segmentation (OVS). OVS refers to the challenging task of segmenting images based on a list of arbitrary text queries. Contrary to traditional segmentation methods, OVS is not limited to a predefined set of classes. The field of OVS has seen significant advancements through different approaches. Approaching OVS as a zero-shot task, pioneering work by ZS3Net [1] and SPNet [29] trained custom modules that translate between visual and linguistic embedding spaces and segment by comparing word vectors and local image features. These methods attempt to learn the challenging connection between two very different and independent latent spaces. Consequently, the field of OVS has seen the introduction of methods incorporating Vision-Language Models (VLMs). VLMs pre-train a visual feature encoder with a textual feature encoder on large amounts of image-text pairs, enabling image and text representations to reside in the same feature space. One of such VLMs, Contrastive Language-Image Pre-training (CLIP) [22], has triggered a wave of segmentation approaches that harness pretrained CLIP embeddings for OVS. LSeg [14] proposed a Dense Prediction Transformer to generate per-pixel CLIP embeddings, which are then compared to class embeddings derived from ground-truth segments. More popularly, two-stage approaches such as OpenSeg [10], OPSNet [3], OVSeg [16], ZSSeg[33] and POMP [24] first obtain class-agnostic masks and use CLIP to classify each mask afterwards. Other works, including MaskCLIP [7], FC-CLIP [36], and SAN [34], focus on utilizing intermediate representations from a frozen CLIP encoder. For example, MaskCLIP uses the token features of the ViT-based CLIP encoder and compares local features to class label embeddings instead of pooling them into a global CLIP embedding. The remarkable diversity in the usage of CLIP for OVS extends to works which leverage relationships between image semantics and class labels in CLIP space [5], use CLIP for binary mask prediction [35] or ensure segmentation consistency at different granularity levels using CLIP [28]. Several works have proposed to tackle OVS outside the realm of VLMs using text-to-image diffusion models [11, 32] or custom transformer networks [30, 17, 31, 38]. Most notably, X-Decoder introduces a generalized transformer model capable of performing various vision-language tasks using a single set of parameters. One common and important limitation of the described methods is the need for textual input as a form of guidance by the user. Our work presents a novel framework which enables self-guidance by automatically finding relevant object categories using localized image captioning. Closest and concurrent to our work is the Zero-Guidance Segmentation paradigm [25], in which clustered DINO [2] embeddings are combined with CLIP to achieve zero guidance by the user. Despite the similar setting, there are key differences. While we use the BLIP [15] model for both image clustering and caption generation, they design a complex pipeline consisting of image clustering with DINO, obtaining vision-language embeddings with CLIP and a custom attention-masking module, to then caption these with CLIP-guided GPT-2. This requires converting between three different latent representations throughout the process. In contrast, we achieve superior segmentation quality by enabling self-guidance using any OVS model.
Image Captioning with Image-Text Embeddings. Image Captioning is the task of describing the content of an image using natural language. A number of approaches have been proposed for the task, including ones utilizing VLMs. Approaches such as CLIP-Cap [20] and CLIP-S [4] utilize CLIP embeddings to guide text generation. Other approaches rely on BLIP [15], a VLM in which the task of decoding text from image features was one of the tasks of the pretraining process. This enables BLIP to perform image captioning without additional components. Furthermore, the text decoder is guided by local patch embeddings in a way that enables local captioning based on a specific area. This unique capability, discovered during the development of our method, emerges organically in BLIP despite being trained solely with image-level captions.
Visual Phrase Grounding. Visual phrase grounding aims to connect different entities mentioned in a caption to corresponding image regions [26]. This task resembles OVS, as it aims to find correspondences between text and image regions. However, predicted areas do not have to precisely outline object boundaries and can be overlapping. Visual grounding has been approached in a self-guided manner, in which heatmaps of regions of the image are generated from image-level CLIP embeddings, which are then captioned by the BLIP encoder [27]. One disadvantage of this approach is that the region determined by CLIP can indicate multiple objects at once, hence multiple objects are captioned for that region, preventing pixel-level class-specific predictions. In our work, BLIP is used both as an encoder and decoder, which enables for category-specific regions and captioning.
Figure 3: Method Overview. BLIP encodings are clustered, aligned and denoised before being decoded into nouns. Generated nouns serve as self-guidance to an OVS model, which outputs the final mask. During evaluation, we process the output mask with our custom evaluator LOVE, which maps predicted nouns to dataset class names.
3 Method
The aim of this paper is to perform open vocabulary semantic segmentation without training, finetuning or textual input. To this end, we propose to identify relevant object categories in the image by captioning local regions in the image at different scales (Sect. 3.1), filtering the generated captions into meaningful semantic descriptions (Sect. 3.2) and using these as self-guidance for a segmentor (Sect. 3.3). An overview of our method is depicted in Fig. 3.
3.1 Local Region Captioning
To identify relevant object categories in the image, we employ Bootstrapped Language Image Pretraining (BLIP) [15]. BLIP is a powerful VLM able to perform accurate and detailed image captioning. The BLIP image encoder is designed as a Vision Transformer [8], dividing the image into patches and encoding them into embeddings in which image-text representations reside. Subsequently, the set of embeddings is passed through transformer layers to capture contextual information across the image. Finally, all contextualized embeddings are passed to the decoder to generate text descriptions. Though effective in describing salient objects in the image, captioning models generally fail to describe all elements of a scene in a comprehensive way. This is likely due to the structure of training data, in which image descriptions mostly focus on salient foreground elements. This characteristic of VLM pre-training limits its direct application to downstream tasks such as open-vocabulary segmentation, in which many objects in the image go unnoticed. As a solution to this limitation, we seek to enable BLIP’s text decoder to consider all encoded areas in the image. Therefore, we propose to cluster the encoded BLIP features into semantically meaningful BLIP feature clusters, which are subsequently enhanced to capture the feature representation of real objects more effectively. Finally, each BLIP feature cluster is fed to the text decoder separately. This enables generating a description of each semantically meaningful area in the image, resulting in a more complete description of the image overall.
Clustering. By default, BLIP expects an RGB image $\mathbf{X}_{D}\in\mathbb{R}^{384\times 384\times 3}$. We additionally process the image at a larger resolution $\mathbf{X}_{L}\in\mathbb{R}^{512\times 512\times 3}$, as common datasets often have images with higher resolution. The multi-resolution set of images $\mathbf{X}=\{\mathbf{X}_{D},\mathbf{X}_{L}\}$ is fed to the BLIP encoder to obtain the set of BLIP patch embeddings $\mathbf{B}$ at the two resolutions:
$\mathbf{P}^{\mathbf{X}_{D}}$ =$=\left\{\mathbf{X}^{D}_{ij}\,|\,\mathbf{X}^{D}_{ij}\in\mathbb{R}^{16\times 16\times 3},1\leq i,j\leq 24\right\}$ (1)
$\mathbf{P}^{\mathbf{X}_{L}}$ =$=\left\{\mathbf{X}^{L}_{ij}\,|\,\mathbf{X}^{L}_{ij}\in\mathbb{R}^{16\times 16\times 3},1\leq i,j\leq 32\right\}$ (2)
$\mathbf{B}^{\mathbf{X}_{D}}$ $=T\big{(}f_{\text{MLP}}(\mathbf{P}^{\mathbf{X}_{D}})\big{)}\quad\textbf{B}^{\mathbf{X}_{L}}=T\big{(}f_{\text{MLP}}(\mathbf{P}^{\mathbf{X}_{L}})\big{)}$ (3)
$\hat{\mathbf{B}}^{\mathbf{X}_{R}}_{i}$ $=\mathbf{B}^{\mathbf{X}_{R}}_{i}\oplus\mathbf{z}^{\mathbf{X}_{R}}_{i},\quad\forall i\in\{1,2,\ldots,N\}$ (4)
where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{L}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the set of patches for the default and large resolutions $D$ and $L$ respectively, $f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP) and $T$ is a Transformer encoder, consisting of $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP:
$T_{l}(X)$ $=\textsc{LayerNorm}\big{(}X+\textsc{MHA}(X,X,X)\big{)}$ (5)
$T^{\prime}_{l}(X)$ $=\textsc{LayerNorm}\big{(}T_{l}(X)+f_{\text{MLP}}(T_{l}(X))\big{)}$ (6)
$T(X)$ $=T^{\prime}_{L}\circ T^{\prime}_{L-1}\circ\ldots\circ T^{\prime}_{1}(X).$ (7)
In Eq. (4), we additionally concatenate a sin-cos positional embedding [^1] Sin-cos positional embeddings are obtained using the positional-encodings Python library. $\mathbf{z}_{i}\in\mathbb{R}^{256}$ to each patch $i$ to encode spatial information, where $N$ is the total number of patches for each resolution $R$. Next, we cluster the patches in $\hat{\mathbf{B}}$ for each resolution $R$ using $k$-means clustering [19]:
[…]
where $C_{k}$ is the set of $k$ clusters and $\mu_{i}$ are the cluster centroids. Running the clustering procedure with two unique values for $k$ on two different image resolutions results in four different cluster assignments.
Cross-clustering Consistency.
Each run of $k$-means clustering labels its clusters independently from others, yielding a correspondence issue between clusters across runs. To resolve this, we relabel the cluster indices to a common reference frame with the following steps:
  1. 1.
    We pick the set of clusters with the most clusters after $k$-means as a reference set $S$. This is not necessarily the set with highest initial k, because some clusters can become empty during the $k$-means iterations. The reference set determines the indices used for all other sets of clusters $\mathcal{C}$, each with its number of clusters denoted by $|C_{i}|$:
    $S=\underset{C_{i}\in\mathcal{C}}{\mathrm{argmax}}\,|C_{i}|$ (9)
  2. 2.
    Sets of clusters are aligned to the reference set using Hungarian matching [13]. Given the reference set $S$ and another set of clusters $C$, we calculate pairwise IoU (Intersection over Union) between the clusters from $R$ and $C$. Then, each cluster from $C$ is assigned a new index, matching the cluster with the highest IoU from $S$:
    $\text{For each cluster }c_{j}\in C,\quad j\in\{1,\ldots,|C|\}:$
    Assign $\text{Assign }c_{j}\text{ to }r_{i}\text{ where }i=\!\underset{i\in\{1,\ldots,|S|\}}{\mathrm{argmax}}\,\text{IoU}(c_{j},r_{i})$ (10)
  3. 3.
    With the labeled sets of clusters, we assign a probability distribution over the clusters to each image patch. For a given patch $p$, let $L(p)=\{L_{1}(p),L_{2}(p),\ldots,L_{m}(p)\}$ be the set of labels assigned to $p$ by the $m$ different sets of clusters. The probability $P(n|p)$ of $p$ being assigned to a particular cluster $n$ is defined as the relative frequency of $n$ among labels $L(p)$:
    $P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }L_{i}(p)=n\\ 0,&\text{\small else}\end{cases}$ (11)
The predictions of a single $k$-means predictor tend to contain high levels of noise, likely due to the high dimensionality of the embeddings. Our method can be seen as an ensemble, reducing the variance present in each individual predictor. The areas that are consistently clustered together by various predictors are likely to be semantically connected. In addition, our method enables for a flexible number of output clusters. Some of the initial clusters can disappear if they are not well-supported by multiple predictors. This property is highly desired due to the variety of input images and the number of objects in them.
Cluster Denoising. To further improve the locality and semantic meaningfulness of clustered feature representations, we apply a Conditional Random Field (CRF) [^2] We use the crfseg Python library. [12] and majority filter. CRF is a discriminative statistical method that is used to denoise predictions based on local interactions between them. In our case, the predictions are a 2D grid of cluster assignment probabilities of the image patches. The CRF implementation we use is specifically tailored for refining 2D segmentation maps. It uses a mean field approximation with a convolutional approach to iteratively adjust the probability distributions of each image patch’s cluster indices. Key to this process is the use of a Gaussian filter in the pairwise potentials, which ensures spatial smoothness and consistency in the segmentation.
As shown in Fig. 3, the application of the CRF yields clusters which are cleaner and more cohesive than the original aligned $k$-means result. To address any remaining noise in the resulting clusters, we apply a neighborhood majority filter as a final step. For each image patch, we consider the set of patches $\mathcal{N}(i,j)$ in its square neighborhood: $\mathcal{N}(i,j)=\{(i+\delta_{i},j+\delta_{j})\,|\,\delta_{i},\delta_{j}\in\{-1,0,1\}\}$. We calculate the mode value from the cluster indices in that neighborhood and assign it as the new index of the central patch:
$\text{mode}\big{(}\mathcal{N}(i,j)\big{)}=\underset{k\in K}{\mathrm{argmax}}\sum_{m\in\mathcal{N}(i,j)}\mathbbm{1}_{\text{index}(m)=k}$ (12)
This step is applied recursively until convergence or 8 times at most.
Caption Generation. The next step involves turning clustered, denoised and enhanced BLIP embeddings into text. The BLIP text decoder is a Transformer architecture capable of processing unordered sets of embeddings of arbitrary size. We leverage this feature and feed flattened subsets of patch embeddings, each corresponding to a cluster, to the BLIP text decoder. Spatial information is preserved due to the presence of positional embeddings added in the clustering step. With this technique, our method essentially iterates over semantic categories captured by clusters and represented by BLIP embeddings. To the best of our knowledge, we are the first to use the text decoder in this manner, enabling local captioning for which it was never specifically trained. The caption generation is stochastic, with different object namings appearing in the captions depending on random initialization. To obtain a rich, unbiased and diverse set of object names, we regenerate captions with each embedding multiple times in caption generation cycles.
3.2 Caption Filtering
The captions generated by the BLIP text decoder are sentences in natural language. For our task, we are only interested in the class names present in each sentence. To obtain these, we filter the sentence down to relevant class names using noun extraction. Captions are parsed using the spaCy model333www.spacy.io/models/en#en_core_web_sm to provide a part-of-speech label for each word in the caption. Words labeled as nouns are kept and converted into their singular form through lemmatization. We collect all nouns generated by different clusters and cycles into one noun set and remove any duplicates, as well as nouns which do not appear in the WordNet dictionary.
3.3 Self-Guidance
The output of the clustered BLIP embeddings (see Sect. 3.1) is a 32x32 grid with cluster index assignments. Each element of the grid corresponds to an image patch from the original image, which enables us to create segmentation masks - defined as a union of areas covered by image patches with the same index - essentially for free. For example, Fig. 3 shows the clustered output as a 2-dimensional mask, which succeed in masking relevant objects to a certain extent. However, the segmentation outputs are constrained to a 32x32 resolution of the BLIP encoder. Upsampling yields unsatisfactory results, since objects appear oversegmented and boundaries remain unsharp. Another challenge is to associate each mask with one class name. In this cluster representation, masked areas are described with sentences rather than single nouns. When such a sentence contains multiple nouns, it is not trivial which noun to select as the predicted class name for that mask. To perform effective and accurate open-vocabulary semantic segmentation, we leverage Blip-Cluster-Caption’s strength in generating an elaborate set of relevant class names and use this as guidance for a pre-trained OVS model capable of producing high-resolution outputs. BLIP-Cluster-Caption is model-agnostic, therefore any kind of OVS model that expects an image and a set of class labels can be combined with our method. We focus on X-Decoder[38], which is a popular and well-performing OVS model.
3.4 Custom Open-Vocabulary Evaluator
As discussed in Sect. 2, previous works in OVS have mainly focused on a setting where the target class names are provided by the user. Hence, evaluation is possible by having access to the ground truth of those target class names during evaluation. However, in scenarios where the target class names are discovered rather than pre-specified, as in our framework, there may be a lack of direct alignment between the semantics of these categories and the classes used in the annotations. In zero-guidance segmentation [25], the authors have proposed to align class names based on the cosine-similarity in the latent space of Sentence-BERT [23] or CLIP [22]. In our initial experiments, this yielded unsatisfactory results where even in obvious cases, the open-vocabulary classes were linked to the incorrect class. Relations between two class names can be complex, such as synonymity, hyponymy, hypernymy, meronymy or holonymy, which exceed the capabilities of a cosine-similarity criterium in the respective latent spaces. As a possible solution, we propose an LLM-based Open-Vocabulary Evaluator (LOVE). Specifically, we employ the Llama 2 7B Large Language model444www.huggingface.co/meta-llama/Llama-2-7b with the task of mapping predicted open-vocabulary categories to target categories in the dataset. First, Self-SEG returns a segmentation mask with classes from an open vocabulary generated by Blip-Cluster-Caption. Second, LOVE collects all unique predicted open-vocabulary categories in the output and builds a mapper from each predicted category to the most relevant target category. Thirdly, all values in the segmentation mask belonging to each class are updated respectively according to the mapper. Finally, the mIoU is computed with the updated segmentation mask. Pseudocode and prompts for LOVE are provided in App. A.
Figure 4: Segmentation Using VLMs. Example outputs from VOC (top), ADE20K (middle), CityScapes (bottom) obtained with different configurations of our self-guidance framework, from which Self-SEG performs most accurately with automatically detected class names.
4 Experiments
4.1 Experimental Setup
We evaluate our method on three popular semantic segmentation validation datasets: Pascal VOC [9], ADE20K [37] and CityScapes [6] with 20, 847 and 20 classes respectively. These datasets cover a wide range of difficulty and class diversity. For example, ADE20K is challenging due to a large variety of classes with many of them appearing only a few times in the whole dataset. We list dataset statistics in Tab. 3. For qualitative comparison with previous works, we use mean Intersection over Union (mIoU) as the main evaluation metric. As mentioned in Sec. 3.4, class predictions in the segmentation mask are mapped using LOVE before the mIoU is computed. To study the class-agnostic segmentation performance of different VLMs, we use class-agnostic mean Intersection over Union (cmIoU). The computation of cmIoU involves the following steps. First, pairwise IoU values, denoted as $\text{IoU}(g_{i},p_{j})$ are calculated between all ground-truth segments $g_{i}$ and predicted segments $p_{j}$ from a given image. Each $g_{i}$ segment is matched with the predicted segment $p_{\text{max}_{i}}$ that yields the highest IoU value: $p_{\text{max}_{i}}=\arg\max_{j}\text{IoU}(g_{i},p_{j})$. The cmIoU is then determined by calculating the arithmetic mean of these highest IoU values across the dataset, which can be formulated as $\text{cmIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}_{\text{max}_{i}}$ where $N$ is the total number of ground-truth segments in the dataset.
In the Self-SEG model used for our experiments, we use the BLIP model built with ViT-Large backbone and finetuned for image captioning on the COCO dataset in combination with the Focal-T variant of X-Decoder. For both models, we use publicly available pretrained weights and do not perform any additional finetuning. Parameter tuning is performed for the clustering and captioning modules. Parameters of the clustering include image scales encoded by BLIP ($384\times 384$ and $512\times 512$), the $k$ values of $k$-means clustering (2 and 8), the parameters of Gaussian smoothing in the CRF (smoothness weight of 6 and smoothness $\theta$ of 0.8), the number of iterations of majority filtering (8) and the feature dimension size of the positional embeddings (256). As for text generation, we opted for nucleus sampling with a minimum length of 4 tokens and maximum of 25, top P value of 1 and repetition penalty of 100 to ensure as many unique nouns as possible. The aforementioned parameters were determined using Bayesian optimization on the Pascal VOC with the goal of maximizing the mIoU.
Figure 5: Qualitative Results. Self-SEG shows remarkable capability to identify open vocabulary class names and segment them accurately across different datasets (top, middle, bottom row are from VOC, ADE20K and CityScapes respectively).
4.2 Ablations
To investigate the performance of different configurations of our framework, we present two ablation studies. The first ablation study investigates different methods of using a VLM to obtain segmentation masks. In the second one, we examine the effect of different amounts of captioning cycles.
Segmentation with Vision Language Models. A VLM can be leveraged in different ways to obtain segmentation masks. In this ablation study, we compare three methods: 1) noun-filtered captions of a plain BLIP encoder/decoder as textual guidance for X-Decoder, which then produces the final segmentation; 2) BLIP cluster embeddings from our Blip-Cluster-Caption approach used as segmentation masks (see Sect. 3.3); and 3) Self-SEG in which textual guidance is achieved through Blip-Cluster-Caption and then provided to X-Decoder. Due to the difficulty with establishing correspondence between BCC clusters and ground-truth labels, we evaluate using cmIoU. Quantitative and qualitative results are reported in Tab. 1 and Fig. 4 respectively. On all three datasets, Self-SEG performs most effectively. Comparing X-Decoder + BLIP to Self-SEG qualitatively, one can observe that the former suffers from false positives and false negatives simultaneously, which in turn serves as an explanation for the performance gap in cmIoU. For example, in Pascal VOC, it incorrectly masks the area around the bus at the cost of segmenting people. In ADE20K, it ignores large parts of the area which could appear to BLIP as insignificant. In our framework, the BLIP-Cluster-Caption ensures that smaller, less salient objects are effectively captioned. As expected, intermediate masks of BLIP-Cluster-Caption do not segment objects sufficiently accurate due to the low resolution of the embedding space. Interestingly, it is still capable of segmenting global areas linked to certain objects in Pascal VOC and ADE20K, while failing completely for CityScapes, likely due to the higher instance count.
Approach VOC [9] ADE20K [37] CityScapes [6]
X-Decoder [38] + BLIP [15] 38.0 26.7 29.2
Blip-Cluster-Caption 16.3 11.3 0.85
Self-SEG 41.9 29.2 35.8
Table 1: Ablations on Segmentation with VLMs. The cmIoU values obtained by different variants of our method. Self-SEG, which combines BCC with X-Decoder segmentation, achieves the best performance on all three datasets.
Captioning Cycles. One important feature of Self-SEG is that it can repeat the process of captioning from BLIP clusters in order to obtain a more elaborate set of nouns. This is particularly useful in situations where objects are initially missed or captioned with a suboptimal semantic description. We analyze the effect of captioning once or repeating the process 5, 10, 15, 20 or 25 times while measuring the mIoU performance. Our results, shown in Tab. 3, indicate that the optimal number of captioning cycles strongly depends on the dataset used. In particular, this number increases with the average amount of instances and unique classes per image (indicated with $\overline{I}$ and $\overline{C}$ in Tab. 3 respectively). For Pascal VOC, a dataset which contains fewer instances and unique classes per image, a single cycle is sufficient for Blip-Cluster-Caption to determine the relevant class names. It should be noted, however, that the average number of instances and classes per image are statistics obtained from the annotations of the dataset. In Pascal VOC, many objects that exist in the images are not annotated and are therefore considered background when evaluated. Hence, the actual optimal number of captioning cycles might be higher. In contrast to Pascal VOC, ADE20K, which contains on average 16 more instances and 7 unique classes per image, requires ten captioning cycles to obtain optimal performance. In the case of CityScapes, images are more complex with 34.3 instances and 17 unique classes on average. Here, twenty cycles of Blip-Cluster-Caption enable the model to obtain an mIoU score of 41.1, whereas Self-SEG would obtain 27.4 mIoU if it had used a single cycle. These outcomes indicate that increasing the number of captioning cycles with higher counts of instances and unique classes per image enable identifying more relevant objects, which in turn benefits the final segmentation output.
Method Auto-Classes VOC [9] ADE20K [37] CityScapes [6]
ZSSeg [33] 83.5 7.0 34.5
FC-CLIP [36] 95.4 14.8 56.2
X-Decoder [38] 96.2 6.4 50.8
OpenSeg [10] 72.2 8.8 -
ODISE [32] 82.7 11.0 -
OVSeg [16] 94.5 9.0 -
SAN [34] 94.6 12.4 -
CatSeg [5] 97.2 13.3 -
LSeg [14] 47.4 - -
OVSegmentor [31] 53.8 - -
SimSeg [35] 57.4 - -
OVDiff [11] 69.0 - -
ZeroSeg [25] 20.1 - -
Self-SEG (Ours) 41.6 6.4 41.1
Table 2: Quantitative Results. Self-SEG performs competitive with existing OVS methods which require class names, while achieving superior performance on the zero label setting.
4.3 Quantitative Analysis
We compare our method against existing OVS methods that require class names to be given, as well as with Zero-Guidance Segmentation (referred to as ZeroSeg [25]). Tab. 2 shows the results. Without any specification of class names, Self-SEG achieves 43%, 43% and 73% of the best performing OVS methods on Pascal VOC, ADE20K and CityScapes respectively. In the ADE20K dataset, Self-SEG is able to remarkably match the performance of X-Decoder [38] without any access to the ground truth class names, while on CityScapes it even outperforms ZSSeg [33] by 6.6 mIoU. It is noteworthy that on CityScapes, our method achieves 81% of X-Decoder’s original performance with given class names. Comparing to ZeroSeg, Self-SEG achieves superior performance on Pascal VOC (41.6 over 20.1 mIoU). Furthermore, we set a new state-of-the-art on ADE20K and CityScapes under the auto-class generation task setting. This outcome underscores the efficacy of our method in dealing with complex scenes that have an open-vocabulary nature, such as ADE20K with its large number of rare classes. While other models have to decide between 847 possibilities, Self-SEG suggests a smaller yet relevant set of class names. Our method deals well with high number of instances and can successfully identify various non-salient objects. A comparison with the findings from Tab. 3 reveals that while Blip-Cluster-Caption struggles with providing high-quality masks, it is able to identify the relevant class names, which can subsequently be segmented accurately using an OVS model. This demonstrates that the integration of Blip-Cluster-Caption and OVS harnesses the strengths of both, leading to enhanced self-guided segmentation performance.
                          Number of Captioning Cycles
Dataset $\overline{I}$ $\overline{C}$ $C$ 1 5 10 15 20 25
Pascal VOC [9] 3.5 3.5 20 41.6 37.9 37.4 35.6 35.1 35.6
ADE20K [37] 19.5 10.5 847 5.1 5.9 6.4 5.9 6.1 6.0
CityScapes [6] 34.3 17 20 27.4 35.9 39.7 40.2 41.1 41.0
Table 3: Ablations on Captioning Cycles. The optimal number of captioning cycles strongly depends on the dataset. Datasets with more instances ($\overline{I}$) and classes on average ($\overline{C}$) benefit from multiple cycles, while captioning once for few instances suffices.
4.4 Qualitative Analysis
Fig. 5 and App. B show qualitative results on the three datasets. Self-SEG is able to exhaustively identify relevant object classes present in the image while segmenting them accurately. Its ability to caption with specific class names such as hawk or treadmill showcases the model’s capabilities for OVS in the wild.
5 Conclusion
In this work we introduced Self-SEG, a novel method which leverages a vision-language model to automatically identify relevant object categories and segment them. In addition, we propose LOVE, a new evaluation method which employs an LLM to map from generated class names to ground-truth labels. Our method performs remarkably, achieving state-of-the-art performance on the zero label setting, while being competitive with open-vocabulary models which require ground-truth labels.
Acknowledgements. This work was supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.
Appendices
Appendix A LLM-based Open-Vocabulary Evaluator
This appendix provides more insight into the pseudocode and text prompts of the LLM Mapper which maps open vocabulary classes to classes in the annotations of the dataset. The LLM Mapper is a key component of our LLM-based Open Vocabulary Evaluation Metric (LOVE) and enables the evaluation of predicted open vocabulary classes.
Pseudocode
The LLM Mapper is tasked with generating a mapping dictionary, which is used to change the predicted class index of each pixel to the corresponding class index in the dataset.
Algorithm 1 LLM Mapper
0:  nouns, dataset_vocabulary, llm_batch_size, prompt_template, LLM
  all_responses, map_dict, skipped $\leftarrow$ initialize()
  for batch in split(nouns, llm_batch_size) do
     dialogs $\leftarrow$ generateDialogs(batch, prompt_template)
     batch_responses $\leftarrow$ LLM(dialogs)
     store(all_responses, (batch_responses, batch))
  end for
  for (batch_response, batch) in all_responses do
     for (response, noun) in (batch_response, batch) do
        answer $\leftarrow$ parseResponse(response)
        common $\leftarrow$ intersect(answer, dataset_vocabulary)
        updateDict(map_dict, noun, common)
        updateSkipped(skipped, noun, common)
     end for
  end for
  for key in map_dict do
     if key is in vocabulary then
        map_dict[key] $\leftarrow$ key
     end if
     if key is in skipped then
        skipped $\leftarrow$ removeItem(skipped, key)
     end if
  end for
  for skipped_key in skipped do
     map_dict[skipped_key] $\leftarrow$ “background”
  end for
  return  map_dict
LLM Prompts
The generateDialogs() function in the LLM Mapper takes a prompt template as one of its inputs. This prompt template is dependent on the dataset and is used to generate dialogs to query the LLM with. For Pascal VOC and CityScapes, we explicitly specify the list of categories, while for ADE20K we rely on the LLM’s knowledge of the dataset. Since ADE20K has many classes, dialogues get very long otherwise. This causes the LLM to run into memory issues when queried with dialogues in batches. We specify the prompt templates for each dataset as follows:
Pascal VOC and CityScapes
To which class in the list <dataset> is
<name> exclusively most similar to? If
<name> is not similar to any class in
the list or if the term describes stuff
instead of things, answer with
’background’. Reply in single quotation
marks with the class name that is part
of the list <dataset> and do not link
it to any other class name which is not
part  of the given list or ’background’.
where <noun> is the noun to map to the vocabulary and <dataset> is the list of classes of Pascal VOC or CityScapes.
ADE20K
To which class in the ADE20K dataset
exclusively most similar to? If <noun>
is not similar to any class in ADE20K
or if the term describes stuff instead
of things, answer with ’background’.
Reply in single quotation marks with
the class name that is part of ADE20K
and do not link it to any other class
name which is not part of the given
list or ’background’.
Appendix B Qualitative Results
This appendix provides additional qualitative results of Self-Seg. With individual sample outputs of the model, we show and discuss both its abilities and limitations.
Figure 6: Self-Seg is able to accurately classify and segment classes not present in the dataset (here Pascal VOC), such as the ox.
Figure 7: Interesting properties of objects are found with Blip-Cluster-Caption, such as reclining by the cat in this image.
Figure 8: Generated nouns are occasionally hyponyms or hypernyms of each other, causing the classes to “compete” for the final segmentation. Future work could possibly address this by checking for such relations between nouns and correcting for them.
Figure 9: Self-guidance occasionally yields a false positive noun, here man, which is then incorrectly used and segmented by X-Decoder.
Figure 10: The generated nouns are occasionally too generic (here animal and afar). The predicted mask is accurate nevertheless.
Figure 11: At times, an object or area is classified with one of its properties rather than the class name itself.
Figure 12: ADE20K has many classes, but badminton was not one of them. Despite being very small, it is still found and segmented.
Figure 13: Even with many objects present, Self-Seg maintains good performance in spotting and segmenting the objects.
Figure 14: Self-Seg can identify and segment an activity present in the image. Shown here is skiing with a white segmentation mask around the feet of the skier.
Figure 15: While in CityScapes all car types are considered equal, Self-Seg often provides more specific descriptions, such as the SUV in this image.
Figure 16: Self-Seg’s freedom in captioning the image with an open-vocabulary understanding of the world becomes evident in this example. Traffic lights in this scene are classified and segmented as signal.
Figure 17: Despite the visual similarities between a moped and motorcycle, Self-Seg correctly considers both. It is able to distinguish between the two vehicles and segment them accurately. A supervised model would not be able to generalize to moped, as it is not one of the annotated classes in CityScapes.

### v2

@@
-Self-Guided Open-Vocabulary Semantic Segmentation
+Auto-Vocabulary Semantic Segmentation
 Abstract
-Vision-Language Models (VLMs) have emerged as promising tools for open-ended image understanding tasks, including open vocabulary segmentation.
-Yet, direct application of such VLMs to segmentation is non-trivial, since VLMs are trained with image-text pairs and naturally lack pixel-level granularity.
-Recent works have made advancements in bridging this gap, often by leveraging the shared image-text space in which the image and a provided text prompt are represented.
-In this paper, we challenge the capabilities of VLMs further and tackle open-vocabulary segmentation without the need for any textual input.
-To this end, we propose a novel Self-Guided Semantic Segmentation (Self-Seg) framework.
-Self-Seg is capable of automatically detecting relevant class names from clustered BLIP embeddings and using these for accurate semantic segmentation.
-In addition, we propose an LLM-based Open-Vocabulary Evaluator (LOVE) to effectively assess predicted open-vocabulary class names.
-We achieve state-of-the-art results on Pascal VOC, ADE20K and CityScapes for open-vocabulary segmentation without given class names, as well as competitive performance with methods where class names are given.
-All code and data will be released.
+Open-ended image understanding tasks gained significant attention from the research community, particularly with the emergence of Vision-Language Models.
+Open-Vocabulary Segmentation (OVS) methods are capable of performing semantic segmentation without relying on a fixed vocabulary, and in some cases, they operate without the need for training or fine-tuning.
+However, OVS methods typically require users to specify the vocabulary based on the task or dataset at hand.
+In this paper, we introduce Auto-Vocabulary Semantic Segmentation (AVS), advancing open-ended image understanding by eliminating the necessity to predefine object categories for segmentation.
+Our approach, AutoSeg, presents a framework that autonomously identifies relevant class names using enhanced BLIP embeddings, which are utilized for segmentation afterwards.
+Given that open-ended object category predictions cannot be directly compared with a fixed ground truth, we develop a Large Language Model-based Auto-Vocabulary Evaluator (LAVE) to efficiently evaluate the automatically generated class names and their corresponding segments.
+Our method sets new benchmarks on datasets such as PASCAL VOC and Context, ADE20K, and Cityscapes for AVS and showcases competitive performance to OVS methods that require specified class names.
+All code is publicly available here.
+Keywords:
+Auto-Vocabulary Semantic Segmentation Open-Vocabulary Semantic Segmentation Semantic Segmentation
 1 Introduction
@@
 While notable advancements have been made, existing semantic segmentation models primarily rely on labeled datasets with predefined categories, limiting their ability to handle unknown classes.
-In contrast, humans possess an open-vocabulary understanding of scenes, recognizing thousands of distinct categories.
+In contrast, humans possess an open-ended understanding of scenes, recognizing thousands of distinct categories.
 Recent studies have focused on the development of segmentation models to meet similar capabilities [24, 20], with models that leverage Vision-Language Models (VLMs) as an emerging category [15, 10, 7, 17, 5, 40, 3, 37, 39].
-Such VLMs learn rich multi-modal features from billions of image-text pairs.
-Figure 1: Self-Guided Open-Vocabulary Semantic Segmentation. Our method is capable of determining and segmenting relevant open vocabulary class names automatically.
+Such VLMs learn rich multi-modal features from large numbers of image-text pairs.
+Figure 1: Different Semantic Segmentation Tasks in Comparison.
+In traditional Semantic Segmentation, an image is segmented into fixed, predefined dataset classes used during annotation.
+In Open-Vocabulary Segmentation, the user specifies which object categories (from the open vocabulary) need to be segmented.
+In Auto-Vocabulary Segmentation, relevant object categories are generated from the image automatically, enabling a true open-ended understanding of scenes without a human in the loop.
+Figure 2: AutoSeg Exemplary Results.
+While established segmentation datasets have a fixed set of annotation categories, our method is able to identify and segment object categories beyond the fixed-set ground truth, such as cub, dachshund or pagoda.
+Images are from the PASCAL VOC [9] and ADE20K [41] datasets.
 Given the scale, training VLMs comes at a extremely high computational cost.
-Because of this, it is common in literature to build upon pre-trained VLMs such as CLIP [22] or BLIP [15].
+Because of this, it is common to build upon pre-trained VLMs such as CLIP [25] or BLIP [16].
 However, applying these VLMs on per-pixel tasks is non-trivial, as they are trained on full images and lack the ability to directly reason over local regions in the image.
@@
 In this paper, we take the capabilities of VLMs one step further and tackle a more challenging task in which the relevant class names are automatically generated.
-We enable self-guided open-vocabulary semantic segmentation, in which the goal is to produce accurate object segmentation and their pixel-level classifications for any class without the need for textual input from the user or predefined class names.
-To tackle this task, we introduce a Self-Guided Semantic Segmentation (Self-SEG) framework.
-Our framework can adapt to any pre-trained segmentation backbone limited by the need for textual input describing the class names to be potentially segmented.
-In order to automate this, we propose BLIP-Cluster-Caption (BCC) which consists of four simple steps: 1) obtain local BLIP embeddings at different image scales; 2) cluster local BLIP embeddings to identify and separate meaningful semantic areas in the image; 3) caption each clustered embedding and extract nouns from the caption; and 4) pass the resulting nouns to a pre-trained segmentation backbone as class names, thereby essentially creating a self-guided network without any additional supervised training.
-Experimental evaluations on Pascal VOC [9], ADE20K [37] and CityScapes [6] showcase the effectiveness of our framework.
-Self-SEG achieves state-of-the-art results on self-guided open-vocabulary semantic segmentation and competitive results with methods where class names are given.
+We enable Auto-Vocabulary Semantic Segmentation (AVS), in which the goal is to produce accurate object segmentation and their pixel-level classifications for any class without the need for textual input from the user, predefined class names, additional data, training or fine-tuning.
+To tackle this task, we introduce the AutoSeg framework.
+Our framework can adapt to any Open-Vocabulary Segmentation (OVS) backbone.
+We propose BLIP-Cluster-Caption (BCC) to generate relevant class names, consisting of four simple steps: 1) obtain local BLIP embeddings at different image scales; 2) cluster local BLIP embeddings to identify and separate meaningful semantic areas in the image; 3) caption each clustered embedding and extract nouns from the caption; and 4) pass the resulting nouns to a pre-trained segmentation backbone as class names, thereby essentially creating a self-guided network.
+Experimental evaluations on PASCAL VOC [9] and Context [23], ADE20K [41] and Cityscapes [6] showcase the effectiveness of AutoSeg, achieving state-of-the-art results on AVS and results competitive to OVS methods which are - unlike AutoSeg- limited by the need for pre-specified class names.
 It is capable of predicting crisp segmentation masks without having any information a priori about which objects to segment (see Fig.
-1).
-For effective evaluation on public datasets, we furthermore propose an LLM-based Open-Vocabulary Evaluator (LOVE), tasked with mapping open-vocabulary predictions to class names in the dataset using a Large Language Model (LLM).
-All code will be publicly released to ensure reproducibility and to facilitate further research.
-In summary, our contributions are as follows: 1) we introduce Self-SEG, a novel framework to automatically determine and segment relevant classes in an image; 2) we propose BLIP-Cluster-Caption, a novel image captioning method capable of generating local and semantically meaningful descriptions from the image obtained from clustered BLIP embeddings; and 3) we propose LOVE, a novel evaluation method for open-vocabulary semantic segmentation which utilizes an LLM.
-Through qualitative and quantitative analyses, we demonstrate the effectiveness of our Self-Guided Semantic Segmentation framework on multiple public datasets.
-Figure 2: Different Semantic Segmentation Tasks in Comparison.
-Left: In traditional semantic segmentation, an image is segmented into fixed, predefined dataset classes used during annotation.
-Middle: In open-vocabulary segmentation, the user specifies which object categories (from the open vocabulary) need to be segmented.
-Right: In self-guided segmentation, relevant object categories are generated from the image automatically, enabling a true open-vocabulary understanding of scenes without a human in the loop.
+2).
+For effective evaluation on public datasets, we furthermore propose a Large Language Model-based Auto-Vocabulary Evaluator (LAVE), tasked with mapping auto-vocabulary predictions to fixed ground truth class names in the dataset.
+All code will be publicly released to ensure reproducibility and to facilitate further research into AVS.
+In summary, our contributions are as follows: 1) we introduce AutoSeg, a novel framework to automatically determine and segment relevant classes in an image; 2) we propose BLIP-Cluster-Caption, a novel image captioning method capable of generating local and semantically meaningful descriptions from the image obtained from enhanced BLIP embeddings; and 3) we propose LAVE, a novel evaluation method for auto-vocabulary semantic segmentation which utilizes an LLM.
+Through qualitative and quantitative analyses, we demonstrate the effectiveness of our framework for Auto-Vocabulary Semantic Segmentation on multiple public datasets.
 2 Related Work
@@
 One of such VLMs, Contrastive Language-Image Pre-training (CLIP) [25], has triggered a wave of segmentation approaches that harness pretrained CLIP embeddings for OVS.
-LSeg [14] proposed a Dense Prediction Transformer to generate per-pixel CLIP embeddings, which are then compared to class embeddings derived from ground-truth segments.
+LSeg [15] employed the Dense Prediction Transformer to generate per-pixel CLIP embeddings, which are then compared to class embeddings derived from ground-truth segments.
 More popularly, two-stage approaches such as OpenSeg [10], OPSNet [3], OVSeg [17], ZSSeg[38] and POMP [27] first obtain class-agnostic masks and use CLIP to classify each mask afterwards.
@@
 This requires converting between three different latent representations throughout the process.
-In contrast, we achieve superior segmentation quality by enabling self-guidance using any OVS model.
+In contrast, we achieve superior segmentation quality by enabling self-guidance using any out-of-the-box OVS model.
 Image Captioning with Image-Text Embeddings.
@@
 Approaches such as CLIP-Cap [22] and CLIP-S [4] utilize CLIP embeddings to guide text generation.
+More recently, the Large Language and Vision Assistant (LLaVA) was introduced with similar capabilities by connecting the visual encoder of CLIP with a language decoder, as well as training on multimodal instruction-following data [18].
 Other approaches rely on BLIP [16], a VLM in which the task of decoding text from image features was one of the tasks of the pretraining process.
@@
 One disadvantage of this approach is that the region determined by CLIP can indicate multiple objects at once, hence multiple objects are captioned for that region, preventing pixel-level class-specific predictions.
-In our work, BLIP is used both as an encoder and decoder, which enables for category-specific regions and captioning.
+In our work, BLIP serves both as encoder and decoder, enabling category-specific regions and captioning.
 Figure 3: Method Overview.
 BLIP encodings are clustered, aligned and denoised before being decoded into nouns.
-Generated nouns serve as self-guidance to an OVS model, which outputs the final mask.
-During evaluation, we process the output mask with our custom evaluator LOVE, which maps predicted nouns to dataset class names.
+Generated nouns serve as self-guidance to a segmentor, which predicts the final mask.
+When evaluating (purple), our custom evaluator LAVE processes the output, mapping predicted nouns to the fixed-vocabulary annotations.
 3 Method
-The aim of this paper is to perform open vocabulary semantic segmentation without training, finetuning or textual input.
-To this end, we propose to identify relevant object categories in the image by captioning local regions in the image at different scales (Sect.
+This work aims to perform auto-vocabulary semantic segmentation without any additional data, training, finetuning or textual input.
+To this end, we propose to identify relevant object categories in the image by captioning local image regions at different scales (Sect.
 3.1), filtering the generated captions into meaningful semantic descriptions (Sect.
-3.2) and using these as self-guidance for a segmentor (Sect.
+3.2) and using the generated classes names as target categories for a segmentor (Sect.
 3.3).
-An overview of our method is depicted in Fig.
+An overview is depicted in Fig.
 3.
@@
 To identify relevant object categories in the image, we employ Bootstrapped Language Image Pretraining (BLIP) [16].
-BLIP is a powerful VLM able to perform accurate and detailed image captioning.
+BLIP is a powerful VLM capable of performing accurate and detailed image captioning.
 The BLIP image encoder is designed as a Vision Transformer [8], dividing the image into patches and encoding them into embeddings in which image-text representations reside.
@@
 Finally, all contextualized embeddings are passed to the decoder to generate text descriptions.
-Though effective in describing salient objects in the image, captioning models generally fail to describe all elements of a scene in a comprehensive way.
-This is likely due to the structure of training data, in which image descriptions mostly focus on salient foreground elements.
-This characteristic of VLM pre-training limits its direct application to downstream tasks such as open-vocabulary segmentation, in which many objects in the image go unnoticed.
-As a solution to this limitation, we seek to enable BLIP’s text decoder to consider all encoded areas in the image.
-Therefore, we propose to cluster the encoded BLIP features into semantically meaningful BLIP feature clusters, which are subsequently enhanced to capture the feature representation of real objects more effectively.
-Finally, each BLIP feature cluster is fed to the text decoder separately.
-This enables generating a description of each semantically meaningful area in the image, resulting in a more complete description of the image overall.
+Though effective in describing salient objects in the image, captioning models like BLIP generally fail to describe all elements of a scene comprehensively.
+This is likely due to the descriptions in the training data, which mostly focus on salient foreground objects.
+This characteristic of VLM pre-training limits its direct application to downstream tasks such as open-ended recognition, in which many objects in the image go unnoticed.
+As a solution to this limitation, we propose BLIP-Cluster-Caption (BCC) to enable the BLIP text decoder to consider all encoded areas in the image.
+BCC clusters the encoded BLIP features into semantically meaningful BLIP feature clusters, which are subsequently enhanced to capture the feature representation of real objects more effectively.
+Finally, each BLIP feature cluster is fed to the BLIP text decoder separately.
+This enables generating a description of each semantically meaningful area in the image, resulting in a more comprehensive and accurate description of the image overall.
 Clustering.
 By default, BLIP expects an RGB image $\mathbf{X}_{D}\in\mathbb{R}^{384\times 384\times 3}$.
-We additionally process the image at a larger resolution $\mathbf{X}_{L}\in\mathbb{R}^{512\times 512\times 3}$, as common datasets often have images with higher resolution.
-The multi-resolution set of images $\mathbf{X}=\{\mathbf{X}_{D},\mathbf{X}_{L}\}$ is fed to the BLIP encoder to obtain the set of BLIP patch embeddings $\mathbf{B}$ at the two resolutions:
+Since common datasets often have images with higher resolution, we additionally process the image at a larger resolution $\mathbf{X}_{H}\in\mathbb{R}^{512\times 512\times 3}$.
+The multi-resolution set of images $\mathbf{X}=\{\mathbf{X}_{D},\mathbf{X}_{H}\}$ is fed to the BLIP encoder to obtain the set of BLIP patch embeddings $\hat{\textbf{B}}=\{\hat{\textbf{B}}^{\mathbf{X}_{D}},\hat{\textbf{B}}^{\mathbf{X}_{H}}\}$ at the two resolutions:
 $\mathbf{P}^{\mathbf{X}_{D}}$ =$=\left\{\mathbf{X}^{D}_{ij}\,|\,\mathbf{X}^{D}_{ij}\in\mathbb{R}^{16\times 16\times 3},1\leq i,j\leq 24\right\}$ (1)
-$\mathbf{P}^{\mathbf{X}_{L}}$ =$=\left\{\mathbf{X}^{L}_{ij}\,|\,\mathbf{X}^{L}_{ij}\in\mathbb{R}^{16\times 16\times 3},1\leq i,j\leq 32\right\}$ (2)
-$\mathbf{B}^{\mathbf{X}_{D}}$ $=T\big{(}f_{\text{MLP}}(\mathbf{P}^{\mathbf{X}_{D}})\big{)}\quad\textbf{B}^{\mathbf{X}_{L}}=T\big{(}f_{\text{MLP}}(\mathbf{P}^{\mathbf{X}_{L}})\big{)}$ (3)
-$\hat{\mathbf{B}}^{\mathbf{X}_{R}}_{i}$ $=\mathbf{B}^{\mathbf{X}_{R}}_{i}\oplus\mathbf{z}^{\mathbf{X}_{R}}_{i},\quad\forall i\in\{1,2,\ldots,N\}$ (4)
-where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{L}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the set of patches for the default and large resolutions $D$ and $L$ respectively, $f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP) and $T$ is a Transformer encoder, consisting of $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP:
-$T_{l}(X)$ $=\textsc{LayerNorm}\big{(}X+\textsc{MHA}(X,X,X)\big{)}$ (5)
-$T^{\prime}_{l}(X)$ $=\textsc{LayerNorm}\big{(}T_{l}(X)+f_{\text{MLP}}(T_{l}(X))\big{)}$ (6)
-$T(X)$ $=T^{\prime}_{L}\circ T^{\prime}_{L-1}\circ\ldots\circ T^{\prime}_{1}(X).$ (7)
+$\mathbf{P}^{\mathbf{X}_{H}}$ =$=\left\{\mathbf{X}^{H}_{ij}\,|\,\mathbf{X}^{H}_{ij}\in\mathbb{R}^{16\times 16\times 3},1\leq i,j\leq 32\right\}$ (2)
+$\hat{\mathbf{B}}^{\mathbf{X}_{R}}_{n}$ $=T\big{(}f_{\text{MLP}}(\mathbf{P}_{n}^{\mathbf{X}_{R}})\big{)}\oplus\mathbf{z}^{\mathbf{X}_{R}}_{n},\quad\forall n\in\{1,2,\ldots,N\}$ (3)
+where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{H}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the set of patches in resolutions $D$ and $H$ respectively, $f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP) and $T$ is a Transformer encoder, consisting of $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP, sequentially propagated with the function composition operator $\circ$:
+$T_{l}(X)$ $=X+\textsc{MHA}(\textsc{LayerNorm}(X,X,X)$ (4)
+$\hat{T}_{l}(X)$ $=T_{l}(X)+f_{\text{MLP}}(\textsc{LayerNorm}(T_{l}(X)))$ (5)
+$T(X)$ $=(\hat{T}_{L-1}\circ\hat{T}_{L-2}\circ\ldots\circ\hat{T}_{0})(X).$ (6)
 In Eq.
-(4), we additionally concatenate a sin-cos positional embedding [^1] Sin-cos positional embeddings are obtained using the positional-encodings Python library.
-$\mathbf{z}_{i}\in\mathbb{R}^{256}$ to each patch $i$ to encode spatial information, where $N$ is the total number of patches for each resolution $R$.
+(3), we additionally concatenate a sinusodial positional encoding [^1] Obtained with the positional-encodings Python library.
+[32] $\mathbf{z}_{i}\in\mathbb{R}^{256}$ to each patch $i$ to encode spatial information, where $N$ is the total number of patches for each resolution $R$.
 Next, we cluster the patches in $\hat{\mathbf{B}}$ for each resolution $R$ using $k$-means clustering [21]:
@@
 where $C_{k}$ is the set of $k$ clusters and $\mu_{i}$ are the cluster centroids.
-Running the clustering procedure with two unique values for $k$ on two different image resolutions results in four different cluster assignments.
+Running the clustering procedure with a 2 to 8 range for $k$ on two different image resolutions results in 14 different cluster assignments.
 Cross-clustering Consistency.
-Each run of $k$-means clustering labels its clusters independently from others, yielding a correspondence issue between clusters across runs.
+Each run of $k$-means clustering labels its clusters independently from others, yielding a correspondence mismatch between clusters across runs.
 To resolve this, we relabel the cluster indices to a common reference frame with the following steps:
-  1. 1.
-    We pick the set of clusters with the most clusters after $k$-means as a reference set $S$.
-This is not necessarily the set with highest initial k, because some clusters can become empty during the $k$-means iterations.
+1.
+Select $C$ with the most clusters after $k$-means as a reference set $S$.
+As some clusters end up empty during the $k$-means iterations, this is not always the set with highest initial $k$.
 The reference set determines the indices used for all other sets of clusters $\mathcal{C}$, each with its number of clusters denoted by $|C_{i}|$:
 $S=\underset{C_{i}\in\mathcal{C}}{\mathrm{argmax}}\,|C_{i}|$ (8)
-  2. 2.
+2.
 Sets of clusters are aligned to the reference set using Hungarian matching [14].
-Given the reference set $S$ and another set of clusters $C$, we calculate pairwise IoU (Intersection over Union) between the clusters from $R$ and $C$.
+We calculate pairwise Intersection over Union (IoU) between the clusters from $S$ and $C$.
 Then, each cluster from $C$ is assigned a new index, matching the cluster with the highest IoU from $S$:
 $\text{For each cluster }c_{j}\in C,\quad j\in\{1,\ldots,|C|\}:$
-    Assign $\text{Assign }c_{j}\text{ to }r_{i}\text{ where }i=\!\underset{i\in\{1,\ldots,|S|\}}{\mathrm{argmax}}\,\text{IoU}(c_{j},r_{i})$ (10)
-  3. 3.
-    With the labeled sets of clusters, we assign a probability distribution over the clusters to each image patch.
-For a given patch $p$, let $L(p)=\{L_{1}(p),L_{2}(p),\ldots,L_{m}(p)\}$ be the set of labels assigned to $p$ by the $m$ different sets of clusters.
-The probability $P(n|p)$ of $p$ being assigned to a particular cluster $n$ is defined as the relative frequency of $n$ among labels $L(p)$:
-    $P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }L_{i}(p)=n\\ 0,&\text{\small else}\end{cases}$ (11)
+Assign index $\text{Assign index }i\text{ to }c_{j}\text{ where }i=\!\underset{i\in\{1,\ldots,|S|\}}{\mathrm{argmax}}\,\text{IoU}(c_{j},s_{i})$ (9)
+3.
+With the labeled sets of clusters, a probability distribution over the clusters is assigned to each image patch.
+For a given patch $p$, let $\mathbf{L}(p)=\{\mathbf{L}_{1}(p),\mathbf{L}_{2}(p),\ldots,\\ \mathbf{L}_{m}(p)\}$ be the set of labels assigned to $p$ by the $m$ different sets of clusters.
+The probability $P(n|p)$ of $p$ being assigned to a particular cluster $n$ is defined as the relative frequency of $n$ among labels $\mathbf{L}(p)$:
+$P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }\mathbf{L}_{i}(p)=n\\ 0,&\text{\small else}\end{cases}$ (10)
 The predictions of a single $k$-means predictor tend to contain high levels of noise, likely due to the high dimensionality of the embeddings.
@@
 In our case, the predictions are a 2D grid of cluster assignment probabilities of the image patches.
-The CRF implementation we use is specifically tailored for refining 2D segmentation maps.
-It uses a mean field approximation with a convolutional approach to iteratively adjust the probability distributions of each image patch’s cluster indices.
+Our implementation is specifically tailored for refining 2D segmentation map, using a mean field approximation with a convolutional approach to iteratively adjust the probability distributions of each image patch’s cluster indices.
 Key to this process is the use of a Gaussian filter in the pairwise potentials, which ensures spatial smoothness and consistency in the segmentation.
-As shown in Fig.
-3, the application of the CRF yields clusters which are cleaner and more cohesive than the original aligned $k$-means result.
-To address any remaining noise in the resulting clusters, we apply a neighborhood majority filter as a final step.
+The application of the CRF yields embeddings which are less noisy and more cohesive than the original aligned $k$-means result.
+To address remaining noise in the embeddings, a neighborhood majority filter is applied as a final step.
 For each image patch, we consider the set of patches $\mathcal{N}(i,j)$ in its square neighborhood: $\mathcal{N}(i,j)=\{(i+\delta_{i},j+\delta_{j})\,|\,\delta_{i},\delta_{j}\in\{-1,0,1\}\}$.
-We calculate the mode value from the cluster indices in that neighborhood and assign it as the new index of the central patch:
+The mode value from the cluster indices in that neighborhood is calculated and assigned as the new index of the central patch:
 $\text{mode}\big{(}\mathcal{N}(i,j)\big{)}=\underset{k\in K}{\mathrm{argmax}}\sum_{m\in\mathcal{N}(i,j)}\mathbbm{1}_{\text{index}(m)=k}$ (11)
-This step is applied recursively until convergence or 8 times at most.
-Caption Generation.
+This step is applied recursively until convergence or 8 times at most. In the supplementary material, we visualize the effect of each step on the embeddings.
+Captioning.
 The next step involves turning clustered, denoised and enhanced BLIP embeddings into text.
@@
 Spatial information is preserved due to the presence of positional embeddings added in the clustering step.
-With this technique, our method essentially iterates over semantic categories captured by clusters and represented by BLIP embeddings.
+With this technique, our method essentially infers semantic categories captured by clusters and represented by BLIP embeddings.
 To the best of our knowledge, we are the first to use the text decoder in this manner, enabling local captioning for which it was never specifically trained.
-The caption generation is stochastic, with different object namings appearing in the captions depending on random initialization.
+The caption generation is stochastic, with different object namings appearing in the captions depending on initialization.
 To obtain a rich, unbiased and diverse set of object names, we regenerate captions with each embedding multiple times in caption generation cycles.
@@
 For example, Fig.
-3 shows the clustered output as a 2-dimensional mask, which succeed in masking relevant objects to a certain extent.
+3 shows the clustered output as a 2-dimensional mask, which succeeds in masking relevant objects to a certain extent.
 However, the segmentation outputs are constrained to a 32x32 resolution of the BLIP encoder.
@@
 When such a sentence contains multiple nouns, it is not trivial which noun to select as the predicted class name for that mask.
-To perform effective and accurate open-vocabulary semantic segmentation, we leverage Blip-Cluster-Caption’s strength in generating an elaborate set of relevant class names and use this as guidance for a pre-trained OVS model capable of producing high-resolution outputs.
-BLIP-Cluster-Caption is model-agnostic, therefore any kind of OVS model that expects an image and a set of class labels can be combined with our method.
+To perform effective and accurate auto-vocabulary semantic segmentation, we leverage BCC’s strength in generating an elaborate set of relevant class names and use this as textual guidance for a pre-trained OVS model capable of producing high-resolution outputs.
+BCC is model-agnostic, therefore any kind of OVS model that expects an image and a set of class labels can be combined with our method.
 We focus on X-Decoder[42], which is a popular and well-performing OVS model.
@@
 In zero-guidance segmentation [28], the authors have proposed to align class names based on the cosine-similarity in the latent space of Sentence-BERT [26] or CLIP [25].
-In our initial experiments, this yielded unsatisfactory results where even in obvious cases, the open-vocabulary classes were linked to the incorrect class.
+In our initial experiments, this yielded unsatisfactory results where, even in obvious cases, the open-vocabulary classes were linked to the incorrect class (e.g.
+taxi to road instead of car).
+This led to a decline in performance as measured by segmentation metrics, yet the method showed promising results from a qualitative perspective.
 Relations between two class names can be complex, such as synonymity, hyponymy, hypernymy, meronymy or holonymy, which exceed the capabilities of a cosine-similarity criterium in the respective latent spaces.
-As a possible solution, we propose an LLM-based Open-Vocabulary Evaluator (LOVE).
-Specifically, we employ the Llama 2 7B Large Language model444www.huggingface.co/meta-llama/Llama-2-7b with the task of mapping predicted open-vocabulary categories to target categories in the dataset.
-First, Self-SEG returns a segmentation mask with classes from an open vocabulary generated by Blip-Cluster-Caption.
-Second, LOVE collects all unique predicted open-vocabulary categories in the output and builds a mapper from each predicted category to the most relevant target category.
-Thirdly, all values in the segmentation mask belonging to each class are updated respectively according to the mapper.
-Finally, the mIoU is computed with the updated segmentation mask.
-Pseudocode and prompts for LOVE are provided in App.
-A.
-Figure 4: Segmentation Using VLMs.
-Example outputs from VOC (top), ADE20K (middle), CityScapes (bottom) obtained with different configurations of our self-guidance framework, from which Self-SEG performs most accurately with automatically detected class names.
+As a possible solution, we propose an LLM-based Auto-Vocabulary Evaluator (LAVE).
+Specifically, we employ the Llama 2 7B Large Language model444www.huggingface.co/meta-llama/Llama-2-7b with the task of mapping predicted auto-vocabulary categories to target categories in the dataset.
+First, AutoSeg returns a segmentation mask with classes from an auto-vocabulary generated by BCC.
+Second, LAVE collects all uniquely predicted auto-vocabulary categories in the output and builds a mapper from each predicted category to the most relevant or similar target category.
+Thirdly, all pixel values in the segmentation mask belonging to each class are updated respectively according to the mapper.
+Finally, the mean Intersection over Union (mIoU) is computed with the updated segmentation mask.
+Pseudocode and prompts for LAVE are provided in the supplementary material.
+Figure 4: Segmentation with VLMs.
+Example outputs on PASCAL VOC/Context (top), ADE20K (middle) and Cityscapes (bottom) by feeding plain BLIP embeddings to X-Decoder, directly using BCC embeddings as masks and AutoSeg.
+Notably, our method segments images in the most comprehensive and accurate manner.
 4 Experiments
 4.1 Experimental Setup
-We evaluate our method on three popular semantic segmentation validation datasets: Pascal VOC [9], ADE20K [37] and CityScapes [6] with 20, 847 and 20 classes respectively.
+We evaluate our method on three popular semantic segmentation validation datasets: PASCAL VOC [9] and Context [23], ADE20K [41] and Cityscapes [6] with 20, 459, 847 and 20 classes respectively.
 These datasets cover a wide range of difficulty and class diversity.
-For example, ADE20K is challenging due to a large variety of classes with many of them appearing only a few times in the whole dataset.
+For example, ADE20K is challenging with its many and infrequently appearing classes, while Cityscapes contains many object instances per image.
 We list dataset statistics in Tab.
-3.
-For qualitative comparison with previous works, we use mean Intersection over Union (mIoU) as the main evaluation metric.
+2.
+Evaluation.
+For quantitative comparison with previous works, we use mIoU as the main metric.
 As mentioned in Sec.
-3.4, class predictions in the segmentation mask are mapped using LOVE before the mIoU is computed.
-To study the class-agnostic segmentation performance of different VLMs, we use class-agnostic mean Intersection over Union (cmIoU).
+3.4, class predictions in the segmentation mask are mapped using LAVE before the mIoU is computed.
+Furthermore, to study the class-agnostic segmentation performance of different VLMs, we report class-agnostic mean Intersection over Union (cmIoU).
 The computation of cmIoU involves the following steps.
@@
 The cmIoU is then determined by calculating the arithmetic mean of these highest IoU values across the dataset, which can be formulated as $\text{cmIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}_{\text{max}_{i}}$ where $N$ is the total number of ground-truth segments in the dataset.
-In the Self-SEG model used for our experiments, we use the BLIP model built with ViT-Large backbone and finetuned for image captioning on the COCO dataset in combination with the Focal-T variant of X-Decoder.
+Table 1: Ablations on Segmentation with VLMs.
+Class-agnostic mIoU values for different variants of our method.
+AutoSeg, which combines BCC with X-Decoder segmentation, achieves the best performance on all four datasets.
+Approach VOC [9] PC-459 [23] ADE20K [41] Cityscapes [6]
+X-Decoder [42] + BLIP [16] 38.0 35.3 26.7 29.2
+Blip-Cluster-Caption 16.3 16.3 11.3 0.85
+AutoSeg 71.8 47.7 29.2 35.8
+Implementation details.
+For our experiments, we use the BLIP model built with ViT-Large backbone and finetuned for image captioning on the COCO dataset in combination with the Focal-L variant of X-Decoder.
 For both models, we use publicly available pretrained weights and do not perform any additional finetuning.
 Parameter tuning is performed for the clustering and captioning modules.
-Parameters of the clustering include image scales encoded by BLIP ($384\times 384$ and $512\times 512$), the $k$ values of $k$-means clustering (2 and 8), the parameters of Gaussian smoothing in the CRF (smoothness weight of 6 and smoothness $\theta$ of 0.8), the number of iterations of majority filtering (8) and the feature dimension size of the positional embeddings (256).
+Parameters of the clustering include image scales encoded by BLIP ($384\times 384$ and $512\times 512$), the $k$ values of $k$-means clustering (2 to 8), the parameters of Gaussian smoothing in the CRF (smoothness weight of 6 and smoothness $\theta$ of 0.8), the number of iterations of majority filtering (8) and the feature dimension size of the positional embeddings (256).
 As for text generation, we opted for nucleus sampling with a minimum length of 4 tokens and maximum of 25, top P value of 1 and repetition penalty of 100 to ensure as many unique nouns as possible.
-The aforementioned parameters were determined using Bayesian optimization on the Pascal VOC with the goal of maximizing the mIoU.
-Figure 5: Qualitative Results.
-Self-SEG shows remarkable capability to identify open vocabulary class names and segment them accurately across different datasets (top, middle, bottom row are from VOC, ADE20K and CityScapes respectively).
+These parameters were determined using Bayesian optimization on VOC with the goal of maximizing the mIoU.
 4.2 Ablations
@@
 The first ablation study investigates different methods of using a VLM to obtain segmentation masks.
-In the second one, we examine the effect of different amounts of captioning cycles.
+In the second, we examine the effect of different amounts of captioning cycles.
 Segmentation with Vision Language Models.
 A VLM can be leveraged in different ways to obtain segmentation masks.
-In this ablation study, we compare three methods: 1) noun-filtered captions of a plain BLIP encoder/decoder as textual guidance for X-Decoder, which then produces the final segmentation; 2) BLIP cluster embeddings from our Blip-Cluster-Caption approach used as segmentation masks (see Sect.
-3.3); and 3) Self-SEG in which textual guidance is achieved through Blip-Cluster-Caption and then provided to X-Decoder.
-Due to the difficulty with establishing correspondence between BCC clusters and ground-truth labels, we evaluate using cmIoU.
-Quantitative and qualitative results are reported in Tab.
-1 and Fig.
-4 respectively.
-On all three datasets, Self-SEG performs most effectively.
-Comparing X-Decoder + BLIP to Self-SEG qualitatively, one can observe that the former suffers from false positives and false negatives simultaneously, which in turn serves as an explanation for the performance gap in cmIoU.
-For example, in Pascal VOC, it incorrectly masks the area around the bus at the cost of segmenting people.
-In ADE20K, it ignores large parts of the area which could appear to BLIP as insignificant.
-In our framework, the BLIP-Cluster-Caption ensures that smaller, less salient objects are effectively captioned.
-As expected, intermediate masks of BLIP-Cluster-Caption do not segment objects sufficiently accurate due to the low resolution of the embedding space.
-Interestingly, it is still capable of segmenting global areas linked to certain objects in Pascal VOC and ADE20K, while failing completely for CityScapes, likely due to the higher instance count.
-Approach VOC [9] ADE20K [37] CityScapes [6]
-X-Decoder [38] + BLIP [15] 38.0 26.7 29.2
-Blip-Cluster-Caption 16.3 11.3 0.85
-Self-SEG 41.9 29.2 35.8
-Table 1: Ablations on Segmentation with VLMs.
-The cmIoU values obtained by different variants of our method.
-Self-SEG, which combines BCC with X-Decoder segmentation, achieves the best performance on all three datasets.
+In this ablation study, we study the benefit of using an auto-vocabulary generated with BCC for segmentation as in AutoSeg.
+In order to do so, we compare our method against a scenario where plain BLIP embeddings are used for captioning instead of enhanced ones.
+Furthermore, we also evaluate the extent to which BCC embedding outputs themselves can be used directly as segmentation masks (see Sect.
+3.3) after upsampling.
+Due to the difficulty of establishing correspondence between BCC clusters and ground-truth labels, we evaluate the segmentation quality using cmIoU.
+Quantitative results are reported in Tab.
+1.
+On all four datasets, AutoSeg performs most effectively.
+Comparing X-Decoder + BLIP to AutoSeg qualitatively, one can observe that the former suffers from false positives and false negatives simultaneously, which in turn serves as an explanation for the performance gap in cmIoU.
+For example, in PASCAL, it incorrectly masks the area around the bus at the cost of segmenting people.
+In ADE20K and Cityscapes, it ignores large parts of the area which could appear to BLIP as insignificant.
+Our framework ensures that smaller, less salient objects are effectively captioned.
+As expected, intermediate masks of BCC do not segment objects sufficiently accurate due to the low resolution of the embedding space.
+Interestingly, it is still capable of segmenting global areas linked to certain objects in PASCAL and ADE20K, while failing completely for Cityscapes, likely due to the higher instance count and higher resolution of images.
+Table 2: Ablations on Captioning Cycles.
+The optimal number of captioning cycles varies per dataset.
+On Cityscapes, which has images in the highest resolution, most instances ($\overline{I}$) and classes on average ($\overline{C}$), AutoSeg benefits from 25 cycles, while between 1 and 15 captioning cycles for datasets with fewer instances per image suffices.
+                                 Number of Captioning Cycles
+Dataset $\overline{I}$ $\overline{C}$ $C$ 1 5 10 15 20 25
+PASCAL VOC [9] 3.5 3.5 20 85.3 84.4 87.1 85.7 83.3 83.3
+PASCAL Context [23] 18.9 6.2 459 10.1 9.7 11.2 11.7 9.6 10.8
+ADE20K [41] 19.5 10.5 847 5.9 5.8 5.1 5.2 5.1 5.6
+Cityscapes [6] 34.3 17 20 28.2 28.1 28.4 28.3 27.9 29.5
 Captioning Cycles.
-One important feature of Self-SEG is that it can repeat the process of captioning from BLIP clusters in order to obtain a more elaborate set of nouns.
-This is particularly useful in situations where objects are initially missed or captioned with a suboptimal semantic description.
-We analyze the effect of captioning once or repeating the process 5, 10, 15, 20 or 25 times while measuring the mIoU performance.
-Our results, shown in Tab.
-3, indicate that the optimal number of captioning cycles strongly depends on the dataset used.
-In particular, this number increases with the average amount of instances and unique classes per image (indicated with $\overline{I}$ and $\overline{C}$ in Tab.
-3 respectively).
-For Pascal VOC, a dataset which contains fewer instances and unique classes per image, a single cycle is sufficient for Blip-Cluster-Caption to determine the relevant class names.
-It should be noted, however, that the average number of instances and classes per image are statistics obtained from the annotations of the dataset.
-In Pascal VOC, many objects that exist in the images are not annotated and are therefore considered background when evaluated.
-Hence, the actual optimal number of captioning cycles might be higher.
-In contrast to Pascal VOC, ADE20K, which contains on average 16 more instances and 7 unique classes per image, requires ten captioning cycles to obtain optimal performance.
-In the case of CityScapes, images are more complex with 34.3 instances and 17 unique classes on average.
-Here, twenty cycles of Blip-Cluster-Caption enable the model to obtain an mIoU score of 41.1, whereas Self-SEG would obtain 27.4 mIoU if it had used a single cycle.
-These outcomes indicate that increasing the number of captioning cycles with higher counts of instances and unique classes per image enable identifying more relevant objects, which in turn benefits the final segmentation output.
-Method Auto-Classes VOC [9] ADE20K [37] CityScapes [6]
-ZSSeg [33] 83.5 7.0 34.5
-FC-CLIP [36] 95.4 14.8 56.2
-X-Decoder [38] 96.2 6.4 50.8
-OpenSeg [10] 72.2 8.8 -
-ODISE [32] 82.7 11.0 -
-OVSeg [16] 94.5 9.0 -
-SAN [34] 94.6 12.4 -
-CatSeg [5] 97.2 13.3 -
-LSeg [14] 47.4 - -
-OVSegmentor [31] 53.8 - -
-SimSeg [35] 57.4 - -
-OVDiff [11] 69.0 - -
-ZeroSeg [25] 20.1 - -
-Self-SEG (Ours) 41.6 6.4 41.1
-Table 2: Quantitative Results. Self-SEG performs competitive with existing OVS methods which require class names, while achieving superior performance on the zero label setting.
+A key aspect of our model is its capability to iteratively repeat the captioning process within BCC, enabling it to generate a more exhaustive collection of nouns which accurately describe the scene.
+This approach proves especially beneficial in instances where objects are initially overlooked or described with less precise semantic terms.
+To assess the impact, we explore the effects of single versus multiple captioning iterations — 5, 10, 15, 20, or 25 times — using the mIoU metric.
+The findings, presented in Tab.
+2, reveal that the optimal number of captioning cycles varies significantly with the dataset in question.
+Specifically, we observe that for the Cityscapes dataset, characterized by its high average number of instances and unique classes per image (denoted as $\overline{I}$ and $\overline{C}$ in Tab.
+2, respectively), a greater number of cycles yields the most benefit.
+For datasets with a lower average number of instances and unique classes per image, like VOC or ADE20K, fewer captioning cycles are required for BCC to identify all relevant class names.
+As anticipated, our model requires a greater number of captioning cycles for the Context dataset, which, despite featuring images identical to those in VOC, includes annotations for an additional 439 object categories.
+Remarkably, AutoSeg successfully handles ADE20K, the dataset with the largest number of unique classes in our experiments, using merely a single captioning cycle.
+This outcome underscores our model’s ability in efficiently addressing complex datasets which resemble an open-ended setting.
+Table 3: Quantitative Results (mIoU).
+AutoSeg performs competitive with existing OVS methods which require class names, while achieving superior performance on the zero label setting.
+Results with dashes are not published and unavailable.
+Method Auto-Classes VOC [9] PC-459 [23] ADE20K [41] Cityscapes [6]
+LSeg [15] 47.4 - - -
+OVSegmentor [36] 53.8 20.4 - -
+SimSeg [39] 57.4 26.2 - -
+OVDiff [11] 69.0 31.4 - -
+OpenSeg [10] 72.2 45.9 8.8 -
+OVSeg [17] 94.5 55.7 9.0 -
+ODISE [34] 82.7 57.3 11.0 -
+SAN [37] 94.6 57.7 12.4 -
+CAT-Seg [5] 97.2 62.0 13.3 -
+ZSSeg [38] 83.5 47.7 7.0 34.5
+X-Decoder [42] 96.2 65.1 6.4 50.8
+FC-CLIP [40] 95.4 58.4 14.8 56.2
+ZeroSeg [28] 20.1 11.4 - -
+AutoSeg (Ours) 87.1 11.7 5.9 29.5
 4.3 Quantitative Analysis
@@
 3 shows the results.
-Without any specification of class names, Self-SEG achieves 43%, 43% and 73% of the best performing OVS methods on Pascal VOC, ADE20K and CityScapes respectively.
-In the ADE20K dataset, Self-SEG is able to remarkably match the performance of X-Decoder [38] without any access to the ground truth class names, while on CityScapes it even outperforms ZSSeg [33] by 6.6 mIoU.
-It is noteworthy that on CityScapes, our method achieves 81% of X-Decoder’s original performance with given class names.
-Comparing to ZeroSeg, Self-SEG achieves superior performance on Pascal VOC (41.6 over 20.1 mIoU).
+Without any specification of class names through user input, AutoSeg achieves 86%, 18%, 40% and 52% of the best performing OVS methods on VOC, Context, ADE20K and Cityscapes respectively.
+The difference in performance can be attributed to AutoSeg needing to identify the subjects for segmentation independently, while OVS methods are explicitly instructed on the specific objects to segment.
+Remarkably, however, AutoSeg is able to remarkably surpass the performance many OVS methods on VOC, such as OVDiff [11], OpenSeg [10] or ODISE [34], without any access to the ground truth class names.
+While still leaving room for improvement, our model compares competitively with OVS methods on ADE20K and Cityscapes, two very challenging datasets either high in number of unique classes or instances.
+Comparing to ZeroSeg, AutoSeg achieves superior performance on VOC (87.1 over 20.1 mIoU) and Context (11.7 over 11.4 mIoU).
 Furthermore, we set a new state-of-the-art on ADE20K and Cityscapes under the auto-class generation task setting.
-This outcome underscores the efficacy of our method in dealing with complex scenes that have an open-vocabulary nature, such as ADE20K with its large number of rare classes.
-While other models have to decide between 847 possibilities, Self-SEG suggests a smaller yet relevant set of class names.
-Our method deals well with high number of instances and can successfully identify various non-salient objects.
+This outcome underscores the efficacy of our method in dealing with complex scenes that have an open-ended nature, such as ADE20K with its large number of rare classes.
+While other models have to decide between 847 possibilities, AutoSeg suggests a smaller yet more relevant set of class names.
+Our method deals well with high numbers of instances and can successfully identify various non-salient objects.
 A comparison with the findings from Tab.
-3 reveals that while Blip-Cluster-Caption struggles with providing high-quality masks, it is able to identify the relevant class names, which can subsequently be segmented accurately using an OVS model.
-This demonstrates that the integration of Blip-Cluster-Caption and OVS harnesses the strengths of both, leading to enhanced self-guided segmentation performance.
-                          Number of Captioning Cycles
-Dataset $\overline{I}$ $\overline{C}$ $C$ 1 5 10 15 20 25
-Pascal VOC [9] 3.5 3.5 20 41.6 37.9 37.4 35.6 35.1 35.6
-ADE20K [37] 19.5 10.5 847 5.1 5.9 6.4 5.9 6.1 6.0
-CityScapes [6] 34.3 17 20 27.4 35.9 39.7 40.2 41.1 41.0
-Table 3: Ablations on Captioning Cycles.
-The optimal number of captioning cycles strongly depends on the dataset.
-Datasets with more instances ($\overline{I}$) and classes on average ($\overline{C}$) benefit from multiple cycles, while captioning once for few instances suffices.
+2 reveals that while BCC struggles with providing high-quality masks on its own, it is able to identify the accurate target classes, which can subsequently be segmented with precision.
+Additional experiments are provided in the supplementary material, in which we generate class names with other image captioning methods and compare the segmentation performance of such vocabularies.
+Overall, our results demonstrate that the integration of BCC and OVS harnesses the strengths of both, leading to enhanced open-ended recognition capabilities.
+Figure 5: Qualitative Results.
+AutoSeg shows remarkable capability to identify out-of-vocabulary categories, such as hawk or coke, and segment them accurately across different datasets.
+The first two rows highlight results on from PASCAL VOC/Context, while the bottom two rows are from ADE20K and Cityscapes respectively.
 4.4 Qualitative Analysis
 Fig.
-5 and App.
-B show qualitative results on the three datasets.
-Self-SEG is able to exhaustively identify relevant object classes present in the image while segmenting them accurately.
-Its ability to caption with specific class names such as hawk or treadmill showcases the model’s capabilities for OVS in the wild.
+5 displays qualitative results on the four datasets, where AutoSeg demonstrates its capability to comprehensively detect and accurately segment relevant object categories within the images.
+Remarkably, it predicts segmentation masks for specific class names such as hawk, piano, wildflower, officer or coke - all categories not listed in the annotations of their respective datasets.
+Moreover, in certain instances, AutoSeg successfully captures additional contextual details, such as graduation or exercise, illustrating the model’s capabilities at recognizing objects in scenes in a genuinely open-ended manner.
+Additional results, as well as failure cases, are available in the supplementary material.
 5 Conclusion
-In this work we introduced Self-SEG, a novel method which leverages a vision-language model to automatically identify relevant object categories and segment them.
-In addition, we propose LOVE, a new evaluation method which employs an LLM to map from generated class names to ground-truth labels.
-Our method performs remarkably, achieving state-of-the-art performance on the zero label setting, while being competitive with open-vocabulary models which require ground-truth labels.
+In this work we introduced AutoSeg, a novel method which leverages a vision-language model to automatically identify relevant target classes and segment them - a task we termed Auto-Vocabulary Semantic Segmentation.
+Additionally, we proposed LAVE, a new evaluation method which employs an LLM to map open-ended class names to ground-truth labels.
+AutoSeg demonstrates remarkable open-ended recognition capabilities, achieving state-of-the-art performance on the zero label setting, while being competitive with open-vocabulary segmentation models which require ground-truth label specification.
 Acknowledgements.
 This work was supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.
-Appendices
-Appendix A LLM-based Open-Vocabulary Evaluator
-This appendix provides more insight into the pseudocode and text prompts of the LLM Mapper which maps open vocabulary classes to classes in the annotations of the dataset.
-The LLM Mapper is a key component of our LLM-based Open Vocabulary Evaluation Metric (LOVE) and enables the evaluation of predicted open vocabulary classes.
-Pseudocode
-The LLM Mapper is tasked with generating a mapping dictionary, which is used to change the predicted class index of each pixel to the corresponding class index in the dataset.
+Supplementary Material
+Appendix A. LLM-based Auto-Vocabulary Evaluator
+This appendix provides more insight into the pseudocode and text prompts of the LLM Mapper which maps auto-vocabulary classes to the fixed-vocabulary classes present in the annotations of each respective dataset.
+The LLM Mapper is a key component of our LLM-based Auto-Vocabulary Evaluation Metric (LAVE) and enables the evaluation of predicted auto-vocabulary classes on public benchmarks.
+By doings so, we enable comparison with with open-vocabulary methods, such as in Tab.
+3, which predict a fixed-vocabulary for evaluation.
+LLM Mapper Pseudocode
+The LLM Mapper is tasked with creating a mapping dictionary which maps from auto-vocabulary classes to fixed-vocabulary classes.
+The resulting mapping dictionary can be used during inference to map pixel-level class predictions belonging to an automatically generated - out-of-vocabulary- class to fixed-vocabulary class indices.
+In Alg.
+1, we detail the procedure to create the mapping dictionary.
+LLM Mapping Prompts
+The generateDialogs() function in the LLM Mapper takes a prompt template as one of its inputs.
+This prompt template is dependent on the dataset and is used to generate dialogs to query the LLM with.
+For PASCAL VOC and Cityscapes, we explicitly specify the list of categories given its short length with 20 classes each.
+For PASCAL Context and ADE20K, with its 459 and 849 classes respectively, we rely on the LLM’s knowledge of the datasets.
+Explicitly naming each object class from these datasets causes the LLM to run into memory issues when queried with long dialogues.
+We specify the prompt templates for each respective dataset as follows:
+PASCAL VOC and Cityscapes
+To which class in the list <dataset> is <name> exclusively most
+similar to? If <name> is not similar to any class in the list or
+if the term describes stuff instead of things, answer with
+‘background’. Reply in single quotation marks with the class name
+that is part of the list <dataset> and do not link it to any other
+class name which is not part of the given list or ‘background’.
+where <noun> is the noun to map to the vocabulary and <dataset> is the list of classes of PASCAL VOC or Cityscapes.
+PASCAL Context and ADE20K
+To which class in the <dataset> dataset is <name> exclusively most
+similar to? If <name> is not similar to any class in <dataset> or
+if the term describes stuff instead of things, answer with
+‘background’. Reply in single quotation marks with the class name
+that is part of <dataset> and do not link it to any other class
+name which is not part of the given list or ‘background’.
+where <noun> is the noun to map to the vocabulary and <dataset> is either ‘PASCAL-Context 459’ or ‘ADE20K’.
+Appendix B. Alternative Class Generation Baselines
+In Tab.
+4, we provide additional comparisons with alternative methods to generate class names from images.
+In particular, we compare against an altered version of BLIP to caption the image at an object instance level with segments retrieved from the Segment Anything Model (SAM[12].
+We use segment everything prompting, i.e.
+the model’s object proposal generation functionality.
+Furthermore, we compare against Large Language and Vision Assistant (LLaVa[18], generating classes with the prompt ‘Please name all the elements present in the image as a list of nouns’.
+We compare the mIoU performance on VOC, Context and Cityscapes and find that AutoSeg performs favorably over other approaches that are adjusted to segment with automatically generated classes.
+The performance gap is especially high for datasets with a small vocabulary, such as VOC and Cityscapes, possibly indicating that AutoSeg is better able in generating classes that are relevant for prediction while the other methods generate redundant classes which become false positive predictions.
+In contrast to both SAM and LLaVa, AutoSeg has an additional benefit of being independent of prompting.
+For this reason, our method does not require prompt engineering in order to obtain class names effective for accurate segmentation.
+Table 4: Experiments with Alternative Class Generation Baselines (mIoU).
+Experiments with alternative methods to generate classes showcase that AutoSeg is favorable for AVS across all three datasets, with larger performance gaps on VOC and Cityscapes with a smaller vocabulary size.
+Approach VOC [9] PC-459 [23] Cityscapes [6]
+SAM + BLIP + X-Decoder 41.1 11.3 27.4
+LLAVA + X-Decoder 56.7 11.4 23.4
+AutoSeg 87.1 11.7 29.5
+Appendix C. Visualization of Embeddings in BCC
+In Fig.
+6, we visualize intermediate 32x32 outputs of BCC to better understand what each individual step of Clustering, Alignment or Denoising achieves.
+Firstly, Fig.
+6 reveals that clustering BLIP embeddings results grouping pixels into semantic areas.
+With $k=2$, the image is divided an area which includes the cows (in blue) and another for everything else (green), while for $k=8$, various other objects - such as the pole in purple - get included.
+At this stage, the embeddings appear noisy and inconsistent across all $k$, which is why the combination of all clusters is so important.
+The alignment output shows that certain objects from a higher $k$ embeddings are now included in those of lower $k$ embeddings, i.e.
+the aligned embedding captures a more extensive list of objects present in the image.
+Finally, to further reduce the noise in the embedding, we apply CRF and a majority filter.
+We see that the final embedding consists of three unique semantic areas, which are eventually fed to the BLIP text decoder for captioning.
+Figure 6: Visualization of Intermediate BCC Outputs. Intermediate BCC outputs show how BLIP embeddings are enhanced with each individual step of Clustering, Alignment or Denoising.
+Appendix D. Additional Qualitative Results
+In Fig. 7,  8 and  9, we provide additional qualitative results of AutoSeg and discuss observed capabilities or limitations in the caption of each figure.
+(a) Success Case. AutoSeg can infer detailed class names, such as courtroom, which is non-existent in the extensive vocabulary of the dataset.
+(b) Success Case. The class badminton is not part of ADE20K. Despite taking up a very small area in the image, AutoSeg is still capable of segmenting it.
+(c) Success Case.
+AutoSeg displays advanced scene understanding capabilities.
+Here, it recognizes the out-of-vocabulary class reflection, which requires the model to reason about other auto-generated classes and structural properties in the image.
+(d) Success Case. The rare and out-of-vocabulary class name dinosaur is generated automatically and segmented by AutoSeg.. This is another example of the method’s efficacy on long-tailed classes.
+(e) Failure Case. AutoSeg’s occasionally struggles with indoor scenes from specific buildings, such as a mosque in this image. Our model mistakenly regards it a temple.
+(f) Failure Case.
+Here, AutoSeg predicts the zebra rug as a zebra as a result of noun parsing.
+After BCC has generated captions from the image, both zebra and rug get parsed into individual nouns.
+When fed to the segmentor, zebra eventually ends up being predicted as the more relevant class name.
+Figure 7: Additional Qualitative Results on ADE20K.
+(a) Success Case.
+Despite the visual similarities between a moped and motorcycle, AutoSeg correctly considers both and segments them accurately.
+A supervised model would fail to generalize to moped, as it is out of Cityscapes’ vocabulary.
+(b) Success Case. While annotations in Cityscapes consider all car types to be equal, AutoSeg often generates more specific descriptions, such as the SUV in this image.
+(c) Success Case.
+AutoSeg is able to provide class names which can be crucial for real-life scenarios, such as hazard in this example.
+The standard 20 classes in Cityscapes do not account for such objects descriptions.
+(d) Failure Case.
+Occasionally, generated class names, such as street, dominate the search space of the segmentor and cause it to attend more on that class.
+This may result in oversegmentation and other class names, such as car, to be ignored.
+Figure 8: Additional Qualitative Results on Cityscapes.
+(a) Success Case. AutoSeg is able to accurately classify and segment classes not present in the dataset, such as the ox.
+(b) Failure Case.
+Generated nouns are occasionally hyponyms or hypernyms of each other, causing the classes to “compete" for the final segmentation.
+Future work could possibly address this by checking for such relations between nouns and correcting for them.
+Figure 9: Additional Qualitative Results on PASCAL.
 Algorithm 1 LLM Mapper
@@
   for skipped_key in skipped do
-     map_dict[skipped_key] $\leftarrow$ “background”
+     map_dict[skipped_key] $\leftarrow$ “background"
   end for
   return  map_dict
-LLM Prompts
-The generateDialogs() function in the LLM Mapper takes a prompt template as one of its inputs.
-This prompt template is dependent on the dataset and is used to generate dialogs to query the LLM with.
-For Pascal VOC and CityScapes, we explicitly specify the list of categories, while for ADE20K we rely on the LLM’s knowledge of the dataset.
-Since ADE20K has many classes, dialogues get very long otherwise.
-This causes the LLM to run into memory issues when queried with dialogues in batches.
-We specify the prompt templates for each dataset as follows:
-Pascal VOC and CityScapes
-To which class in the list <dataset> is
-<name> exclusively most similar to? If
-<name> is not similar to any class in
-the list or if the term describes stuff
-instead of things, answer with
-’background’. Reply in single quotation
-marks with the class name that is part
-of the list <dataset> and do not link
-it to any other class name which is not
-part  of the given list or ’background’.
-where <noun> is the noun to map to the vocabulary and <dataset> is the list of classes of Pascal VOC or CityScapes.
-ADE20K
-To which class in the ADE20K dataset
-exclusively most similar to? If <noun>
-is not similar to any class in ADE20K
-or if the term describes stuff instead
-of things, answer with ’background’.
-Reply in single quotation marks with
-the class name that is part of ADE20K
-and do not link it to any other class
-name which is not part of the given
-list or ’background’.
-Appendix B Qualitative Results
-This appendix provides additional qualitative results of Self-Seg. With individual sample outputs of the model, we show and discuss both its abilities and limitations.
-Figure 6: Self-Seg is able to accurately classify and segment classes not present in the dataset (here Pascal VOC), such as the ox.
-Figure 7: Interesting properties of objects are found with Blip-Cluster-Caption, such as reclining by the cat in this image.
-Figure 8: Generated nouns are occasionally hyponyms or hypernyms of each other, causing the classes to “compete” for the final segmentation.
-Future work could possibly address this by checking for such relations between nouns and correcting for them.
-Figure 9: Self-guidance occasionally yields a false positive noun, here man, which is then incorrectly used and segmented by X-Decoder.
-Figure 10: The generated nouns are occasionally too generic (here animal and afar). The predicted mask is accurate nevertheless.
-Figure 11: At times, an object or area is classified with one of its properties rather than the class name itself.
-Figure 12: ADE20K has many classes, but badminton was not one of them. Despite being very small, it is still found and segmented.
-Figure 13: Even with many objects present, Self-Seg maintains good performance in spotting and segmenting the objects.
-Figure 14: Self-Seg can identify and segment an activity present in the image. Shown here is skiing with a white segmentation mask around the feet of the skier.
-Figure 15: While in CityScapes all car types are considered equal, Self-Seg often provides more specific descriptions, such as the SUV in this image.
-Figure 16: Self-Seg’s freedom in captioning the image with an open-vocabulary understanding of the world becomes evident in this example.
-Traffic lights in this scene are classified and segmented as signal.
-Figure 17: Despite the visual similarities between a moped and motorcycle, Self-Seg correctly considers both.
-It is able to distinguish between the two vehicles and segment them accurately.
-A supervised model would not be able to generalize to moped, as it is not one of the annotated classes in CityScapes.

### v3

@@
 Abstract
-Open-ended image understanding tasks gained significant attention from the research community, particularly with the emergence of Vision-Language Models.
-Open-Vocabulary Segmentation (OVS) methods are capable of performing semantic segmentation without relying on a fixed vocabulary, and in some cases, they operate without the need for training or fine-tuning.
-However, OVS methods typically require users to specify the vocabulary based on the task or dataset at hand.
+Open-Vocabulary Segmentation (OVS) methods are capable of performing semantic segmentation without relying on a fixed vocabulary, and in some cases, without training or fine-tuning.
+However, OVS methods typically require a human in the loop to specify the vocabulary based on the task or dataset at hand.
 In this paper, we introduce Auto-Vocabulary Semantic Segmentation (AVS), advancing open-ended image understanding by eliminating the necessity to predefine object categories for segmentation.
-Our approach, AutoSeg, presents a framework that autonomously identifies relevant class names using enhanced BLIP embeddings, which are utilized for segmentation afterwards.
-Given that open-ended object category predictions cannot be directly compared with a fixed ground truth, we develop a Large Language Model-based Auto-Vocabulary Evaluator (LAVE) to efficiently evaluate the automatically generated class names and their corresponding segments.
-Our method sets new benchmarks on datasets such as PASCAL VOC and Context, ADE20K, and Cityscapes for AVS and showcases competitive performance to OVS methods that require specified class names.
-All code is publicly available here.
-Keywords:
-Auto-Vocabulary Semantic Segmentation Open-Vocabulary Semantic Segmentation Semantic Segmentation
+Our approach, AutoSeg, presents a framework that autonomously identifies relevant class names using semantically enhanced BLIP embeddings and segments them afterwards.
+Given that open-ended object category predictions cannot be directly compared with a fixed ground truth, we develop a Large Language Model-based Auto-Vocabulary Evaluator (LAVE) to efficiently evaluate the automatically generated classes and their corresponding segments.
+With AVS, our method sets new benchmarks on datasets PASCAL VOC, Context, ADE20K, and Cityscapes, while showing competitive performance to OVS methods that require specified class names.
+All code will be publicly released.
+Fixed-Set Ground Truth [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+AutoSeg (Ours) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Open-Ended Predictions Out-of-Vocabulary Predictions
+Figure 1: AutoSeg Exemplary Results.
+AutoSeg is readily applicable to unseen images for open-ended segmentation for objects such as mascot and hole, such as the two images on the left.
+Furthermore, where established segmentation datasets have a fixed set of annotation categories, our method is able to identify and segment with more semantically precise object categories beyond the fixed-set ground truth, such as dachshund, bed and pagoda.
+Images are from the Road Anomaly [20], PASCAL [9] and ADE20K [47] datasets.
 1 Introduction
-Semantic segmentation is a computer vision task that involves grouping pixels into coherent regions based on their semantic categories.
-While notable advancements have been made, existing semantic segmentation models primarily rely on labeled datasets with predefined categories, limiting their ability to handle unknown classes.
-In contrast, humans possess an open-ended understanding of scenes, recognizing thousands of distinct categories.
-Recent studies have focused on the development of segmentation models to meet similar capabilities [24, 20], with models that leverage Vision-Language Models (VLMs) as an emerging category [15, 10, 7, 17, 5, 40, 3, 37, 39].
+While humans possess an open-ended understanding of scenes, recognizing thousands of distinct categories, semantic segmentation methods [25] typically rely on a fixed vocabulary of predefined semantic categories.
+They require large human-annotated datasets and have limited capabilities for handling a broad range of classes or unknown objects.
+Recent studies have focused on addressing these limitations [28, 23] with models that leverage Vision-Language Models (VLMs) as an emerging category [17, 10, 7, 19, 5, 46, 3, 43, 45].
 Such VLMs learn rich multi-modal features from large numbers of image-text pairs.
-Figure 1: Different Semantic Segmentation Tasks in Comparison.
-In traditional Semantic Segmentation, an image is segmented into fixed, predefined dataset classes used during annotation.
-In Open-Vocabulary Segmentation, the user specifies which object categories (from the open vocabulary) need to be segmented.
-In Auto-Vocabulary Segmentation, relevant object categories are generated from the image automatically, enabling a true open-ended understanding of scenes without a human in the loop.
-Figure 2: AutoSeg Exemplary Results.
-While established segmentation datasets have a fixed set of annotation categories, our method is able to identify and segment object categories beyond the fixed-set ground truth, such as cub, dachshund or pagoda.
-Images are from the PASCAL VOC [9] and ADE20K [41] datasets.
-Given the scale, training VLMs comes at a extremely high computational cost.
-Because of this, it is common to build upon pre-trained VLMs such as CLIP [25] or BLIP [16].
-However, applying these VLMs on per-pixel tasks is non-trivial, as they are trained on full images and lack the ability to directly reason over local regions in the image.
-Previous attempts have employed a two-stage approach where masked regions are fed to the VLMs to produce CLIP embeddings, which are subsequently correlated with text encodings of text prompts to perform classification [38, 17, 7].
-Other works directly align per-pixel embeddings with text embeddings retrieved from CLIP [15] based on text prompts provided by the user at test time.
-While training these models, the text prompts usually correspond to the ground truth class names of the dataset in question.
-In this paper, we take the capabilities of VLMs one step further and tackle a more challenging task in which the relevant class names are automatically generated.
-We enable Auto-Vocabulary Semantic Segmentation (AVS), in which the goal is to produce accurate object segmentation and their pixel-level classifications for any class without the need for textual input from the user, predefined class names, additional data, training or fine-tuning.
-To tackle this task, we introduce the AutoSeg framework.
-Our framework can adapt to any Open-Vocabulary Segmentation (OVS) backbone.
-We propose BLIP-Cluster-Caption (BCC) to generate relevant class names, consisting of four simple steps: 1) obtain local BLIP embeddings at different image scales; 2) cluster local BLIP embeddings to identify and separate meaningful semantic areas in the image; 3) caption each clustered embedding and extract nouns from the caption; and 4) pass the resulting nouns to a pre-trained segmentation backbone as class names, thereby essentially creating a self-guided network.
-Experimental evaluations on PASCAL VOC [9] and Context [23], ADE20K [41] and Cityscapes [6] showcase the effectiveness of AutoSeg, achieving state-of-the-art results on AVS and results competitive to OVS methods which are - unlike AutoSeg- limited by the need for pre-specified class names.
-It is capable of predicting crisp segmentation masks without having any information a priori about which objects to segment (see Fig.
+Figure 2: Semantic Segmentation Tasks in Comparison.
+In traditional Semantic Segmentation, an image is segmented into fixed, predefined set of classes (fixed vocabulary).
+In Open-Vocabulary Segmentation, the user specifies which object categories (from the open vocabulary) should be segmented: 1) either via a human-provided prompt at runtime, or 2) the OV-method is trained to output the vocabulary of a human-annotated target dataset.
+In contrast, Auto-Vocabulary Segmentation automatically generates relevant object categories directly from the image.
+This enables true open-ended scene understanding without needing human input.
+Due to the immense computational training costs of VLMs, it is common to build upon pre-trained VLMs such as CLIP [29] or BLIP [18].
+However, applying these VLMs on per-pixel tasks to obtain precise locality information is non-trivial, as they are trained on full images and lack the ability to directly reason over local regions in the image.
+This makes them less useful for obtaining precise segmentation boundaries, which is crucial for downstream tasks like robot grasping.
+Moreover, current Open-Vocabulary segmentation (OVS) methods still require a human in the loop, inherently limiting the scalability of these methods (see Fig.
 2).
-For effective evaluation on public datasets, we furthermore propose a Large Language Model-based Auto-Vocabulary Evaluator (LAVE), tasked with mapping auto-vocabulary predictions to fixed ground truth class names in the dataset.
-All code will be publicly released to ensure reproducibility and to facilitate further research into AVS.
-In summary, our contributions are as follows: 1) we introduce AutoSeg, a novel framework to automatically determine and segment relevant classes in an image; 2) we propose BLIP-Cluster-Caption, a novel image captioning method capable of generating local and semantically meaningful descriptions from the image obtained from enhanced BLIP embeddings; and 3) we propose LAVE, a novel evaluation method for auto-vocabulary semantic segmentation which utilizes an LLM.
-Through qualitative and quantitative analyses, we demonstrate the effectiveness of our framework for Auto-Vocabulary Semantic Segmentation on multiple public datasets.
+OVS methods can be separated into two categories: 1) prompt-based methods (e.g.
+[13, 16, 48]) require per-image user input to provide a known vocabulary; and 2) dataset-based methods (e.g.
+[10, 19, 17, 41]) require human-annotated datasets in combination with training in which the known vocabulary is baked into the method as its defined output categorization.
+For both categories, a human is providing a known and fixed vocabulary, which drastically simplifies the segmentation problem as the ground truth vocabulary is always provided, but it also limits the application use-cases when a human needs to be in the loop.
+Imagine a kitchen robot that needs to distinguish a large variety of tools and ingredients when cooking a recipe and requiring precise segmentations for grasping.
+It is undesirable to require a human-provided prompt for every grasp or to re-label a dataset and re-train the segmentation method every time a new recipe contains a new tool or ingredient.
+To eliminate the human in the loop and fully rely on pre-trained foundation models, we propose Auto-Vocabulary Semantic Segmentation (AVS).
+In AVS, the unknown vocabulary is automatically generated and integrated into a semantic segmentation method thus allowing pixel-level classifications for any class without the need for textual input from the user, predefined class names, additional data, training or fine-tuning.
+To address this task, we introduce AutoSeg, a zero-shot method that first identifies image-specific target categories and then predicts masks for them.
+See Fig.
+1 for exemplary results.
+We introduce BBoost for vocabulary generation, which provides a simple yet effective semantic clustering strategy to enhance locality and semantic precision when captioning with BLIP.
+[…]
+In summary, our contributions are as follows: 1) we introduce AutoSeg, a novel framework to automatically determine and segment open-ended classes in an image; 2) we propose BBoost, a novel method for generating image-specific target vocabularies, leveraging text decoding from enhanced BLIP embeddings; and 3) we propose LAVE, a novel evaluation approach for auto-vocabulary semantic segmentation which utilizes an LLM.
+Through qualitative and quantitative analyses, we demonstrate the effectiveness of our framework for AVS on multiple public datasets.
+Figure 3: Method Overview.
+BLIP encodings are clustered, aligned and denoised before being decoded into nouns by BBoost.
+Generated nouns serve as self-guidance to a segmentor, which predicts the final mask.
+When evaluating (purple), our custom evaluator LAVE processes the output, mapping predicted nouns to the fixed-vocabulary annotations.
 2 Related Work
 Open-Vocabulary Segmentation (OVS).
-OVS refers to the challenging task of segmenting images based on a list of arbitrary text queries.
-Contrary to traditional segmentation methods, OVS is not limited to a predefined set of classes.
-The field of OVS has seen significant advancements through different approaches.
-Approaching OVS as a zero-shot task, pioneering work by ZS3Net [1] and SPNet [33] trained custom modules that translate between visual and linguistic embedding spaces and segment by comparing word vectors and local image features.
-These methods attempt to learn the challenging connection between two very different and independent latent spaces.
-Consequently, the field of OVS has seen the introduction of methods incorporating Vision-Language Models (VLMs).
-VLMs pre-train a visual feature encoder with a textual feature encoder on large amounts of image-text pairs, enabling image and text representations to reside in the same feature space.
-One of such VLMs, Contrastive Language-Image Pre-training (CLIP) [25], has triggered a wave of segmentation approaches that harness pretrained CLIP embeddings for OVS.
-LSeg [15] employed the Dense Prediction Transformer to generate per-pixel CLIP embeddings, which are then compared to class embeddings derived from ground-truth segments.
-More popularly, two-stage approaches such as OpenSeg [10], OPSNet [3], OVSeg [17], ZSSeg[38] and POMP [27] first obtain class-agnostic masks and use CLIP to classify each mask afterwards.
-Other works, including MaskCLIP [7], FC-CLIP [40], and SAN [37], focus on utilizing intermediate representations from a frozen CLIP encoder.
-For example, MaskCLIP uses the token features of the ViT-based CLIP encoder and compares local features to class label embeddings instead of pooling them into a global CLIP embedding.
-The remarkable diversity in the usage of CLIP for OVS extends to works which leverage relationships between image semantics and class labels in CLIP space [5], use CLIP for binary mask prediction [39] or ensure segmentation consistency at different granularity levels using CLIP [31].
-Several works have proposed to tackle OVS outside the realm of VLMs using text-to-image diffusion models [11, 34] or custom transformer networks [35, 19, 36, 42].
-Most notably, X-Decoder introduces a generalized transformer model capable of performing various vision-language tasks using a single set of parameters.
-One common and important limitation of the described methods is the need for textual input as a form of guidance by the user.
-Our work presents a novel framework which enables self-guidance by automatically finding relevant object categories using localized image captioning.
-Closest and concurrent to our work is the Zero-Guidance Segmentation paradigm [28], in which clustered DINO [2] embeddings are combined with CLIP to achieve zero guidance by the user.
-Despite the similar setting, there are key differences.
-While we use the BLIP [16] model for both image clustering and caption generation, they design a complex pipeline consisting of image clustering with DINO, obtaining vision-language embeddings with CLIP and a custom attention-masking module, to then caption these with CLIP-guided GPT-2.
-This requires converting between three different latent representations throughout the process.
-In contrast, we achieve superior segmentation quality by enabling self-guidance using any out-of-the-box OVS model.
+OVS tackles the challenging task of segmenting images based on arbitrary text queries rather than a predefined class set.
+Early OVS methods, like ZS3Net [1] and SPNet [38], treated it as a zero-shot task by training modules that align visual and linguistic embeddings, segmenting by comparing word vectors with local image features.
+Recently, Vision-Language Models (VLMs) pre-trained on large image-text dataset, such as Contrastive Language-Image Pre-training (CLIP)[29], have unified image and text features.
+CLIP has advanced the field of OVS significantly, with methods such as LSeg[17], which compares per-pixel embeddings with class embeddings, and two-stage approaches like OpenSeg [10], OPSNet [3], OVSeg [19], ZSSeg [42], and POMP [31], which first obtain class-agnostic masks and then classify each mask with CLIP.
+Other works, such as MaskCLIP [7], FC-CLIP [46], and SAN [43], utilize intermediate CLIP representations.
+Other approaches leverage CLIP to relate image semantics to class labels [5], predict binary masks [45] or ensure segmentation consistency at multiple granularities [35].
+Beyond VLMs, recent models use text-to-image diffusion [12, 41] or transformers [39, 22, 40, 48], with X-Decoder providing a generalized model for various vision-language tasks.One common and significant limitation of the described methods is the need for textual input as a form of guidance by the user.
+Our work presents a novel framework which enables self-guidance by automatically identifying relevant object categories via localized image captioning.
+Closest and concurrent to our work is Zero-Guidance Segmentation [32], in which clustered DINO [2] embeddings are combined with CLIP to achieve zero guidance by the user.
+Despite the similar setting, there are key differences: we use BLIP [18] for both image clustering and caption generation, while they use a complex pipeline involving DINO clustering, CLIP embeddings, a custom attention-masking module, and CLIP-guided GPT-2 for captioning, requiring conversions between three latent representations.
+In contrast, we achieve superior segmentation quality with self-guidance using any out-of-the-box OVS model.
 Image Captioning with Image-Text Embeddings.
@@
 In our work, BLIP serves both as encoder and decoder, enabling category-specific regions and captioning.
-Figure 3: Method Overview.
-BLIP encodings are clustered, aligned and denoised before being decoded into nouns.
-Generated nouns serve as self-guidance to a segmentor, which predicts the final mask.
-When evaluating (purple), our custom evaluator LAVE processes the output, mapping predicted nouns to the fixed-vocabulary annotations.
 3 Method
-This work aims to perform auto-vocabulary semantic segmentation without any additional data, training, finetuning or textual input.
-To this end, we propose to identify relevant object categories in the image by captioning local image regions at different scales (Sect.
-3.1), filtering the generated captions into meaningful semantic descriptions (Sect.
-3.2) and using the generated classes names as target categories for a segmentor (Sect.
-3.3).
+This work aims to perform semantic segmentation without any additional data, training, finetuning or pre-defined target categories.
+To this end, we propose to identify relevant object categories in an image by captioning with locality (Sect.
+3.1), filter the generated captions into a meaningful vocabulary and use the vocabulary as target categories for a segmentor (Sect.
+3.2).
 An overview is depicted in Fig.
@@
 BLIP is a powerful VLM capable of performing accurate and detailed image captioning.
-The BLIP image encoder is designed as a Vision Transformer [8], dividing the image into patches and encoding them into embeddings in which image-text representations reside.
-Subsequently, the set of embeddings is passed through transformer layers to capture contextual information across the image.
-Finally, all contextualized embeddings are passed to the decoder to generate text descriptions.
+Its image encoder is designed as a Vision Transformer [8], dividing the image into patches and encoding them into embeddings in which image-text features reside.
+The set of embeddings is then passed through transformer layers to capture contextual information across the image.
+Finally, all contextualized embeddings are decoded into text descriptions.
 Though effective in describing salient objects in the image, captioning models like BLIP generally fail to describe all elements of a scene comprehensively.
-This is likely due to the descriptions in the training data, which mostly focus on salient foreground objects.
-This characteristic of VLM pre-training limits its direct application to downstream tasks such as open-ended recognition, in which many objects in the image go unnoticed.
-As a solution to this limitation, we propose BLIP-Cluster-Caption (BCC) to enable the BLIP text decoder to consider all encoded areas in the image.
-BCC clusters the encoded BLIP features into semantically meaningful BLIP feature clusters, which are subsequently enhanced to capture the feature representation of real objects more effectively.
-Finally, each BLIP feature cluster is fed to the BLIP text decoder separately.
-This enables generating a description of each semantically meaningful area in the image, resulting in a more comprehensive and accurate description of the image overall.
+This is likely due to the captions in the training data, which mostly focus on salient foreground objects.
+In turn, this limits its direct application to downstream tasks such as open-ended recognition, since many non-salient objects are missed.
+As a solution to this limitation, we propose BBoost for exhaustive, local captioning.
+BBoost clusters the encoded BLIP features into semantically meaningful feature clusters, which are enhanced to capture the object-specific feature representations more effectively.
+Afterwards, each individual feature cluster is fed to a pre-trained BLIP text decoder.
+This enables generating a description of each semantically distinct and meaningful area in the image, resulting in a more comprehensive, specific and accurate description of the image overall.
+Its components are detailed in the next paragraphs.
 Clustering.
 By default, BLIP expects an RGB image $\mathbf{X}_{D}\in\mathbb{R}^{384\times 384\times 3}$.
-Since common datasets often have images with higher resolution, we additionally process the image at a larger resolution $\mathbf{X}_{H}\in\mathbb{R}^{512\times 512\times 3}$.
+Since common datasets often have images with higher resolution, we additionally process the image at a higher resolution $\mathbf{X}_{H}\in\mathbb{R}^{512\times 512\times 3}$.
 The multi-resolution set of images $\mathbf{X}=\{\mathbf{X}_{D},\mathbf{X}_{H}\}$ is fed to the BLIP encoder to obtain the set of BLIP patch embeddings $\hat{\textbf{B}}=\{\hat{\textbf{B}}^{\mathbf{X}_{D}},\hat{\textbf{B}}^{\mathbf{X}_{H}}\}$ at the two resolutions:
@@
 $\mathbf{P}^{\mathbf{X}_{H}}$ =$=\left\{\mathbf{X}^{H}_{ij}\,|\,\mathbf{X}^{H}_{ij}\in\mathbb{R}^{16\times 16\times 3},1\leq i,j\leq 32\right\}$ (2)
-$\hat{\mathbf{B}}^{\mathbf{X}_{R}}_{n}$ $=T\big{(}f_{\text{MLP}}(\mathbf{P}_{n}^{\mathbf{X}_{R}})\big{)}\oplus\mathbf{z}^{\mathbf{X}_{R}}_{n},\quad\forall n\in\{1,2,\ldots,N\}$ (3)
-where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{H}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the set of patches in resolutions $D$ and $H$ respectively, $f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP) and $T$ is a Transformer encoder, consisting of $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP, sequentially propagated with the function composition operator $\circ$:
-$T_{l}(X)$ $=X+\textsc{MHA}(\textsc{LayerNorm}(X,X,X)$ (4)
-$\hat{T}_{l}(X)$ $=T_{l}(X)+f_{\text{MLP}}(\textsc{LayerNorm}(T_{l}(X)))$ (5)
+$\hat{\mathbf{B}}_{n}^{\mathbf{X}_{R}}$ $=T\left(f_{\text{MLP}}\left(\mathbf{P}_{n}^{\mathbf{X}_{R}}\right)\right)\parallel\mathbf{z}_{n}^{\mathbf{X}_{R}},$
+$\quad\forall R\in\{D,H\},\quad n\in\{1,\ldots,N_{R}\}$ (3)
+where $\textbf{P}^{\textbf{X}_{D}}\in\mathbb{R}^{576\times 16\times 16\times 3}$ and $\textbf{P}^{\textbf{X}_{H}}\in\mathbb{R}^{1024\times 16\times 16\times 3}$ denote the sets of patches for resolutions $D$ and $H$ respectively, and $\mathbf{P}_{n}^{\mathbf{X}_{R}}$ is the $n$-th patch.
+$f_{\text{MLP}}(\cdot)$ represents a shared fully connected Multi-Layer Perceptron (MLP), LN is LayerNorm, and $T$ is a Transformer encoder with $L$ alternating layers of Multi-Head Self-Attention (MHA) and an MLP, sequentially propagated:
+$T_{l}(X)$ $=X+\textsc{MHA}\big{(}\textsc{LN}(X),\textsc{LN}(X),\textsc{LN}(X)\big{)}$ (4)
+$\hat{T}_{l}(X)$ $=T_{l}(X)+f_{\text{MLP}}\big{(}\textsc{LN}(T_{l}(X))\big{)}$ (5)
 $T(X)$ $=(\hat{T}_{L-1}\circ\hat{T}_{L-2}\circ\ldots\circ\hat{T}_{0})(X).$ (6)
 In Eq.
-(3), we additionally concatenate a sinusodial positional encoding [^1] Obtained with the positional-encodings Python library.
-[32] $\mathbf{z}_{i}\in\mathbb{R}^{256}$ to each patch $i$ to encode spatial information, where $N$ is the total number of patches for each resolution $R$.
-Next, we cluster the patches in $\hat{\mathbf{B}}$ for each resolution $R$ using $k$-means clustering [21]:
+(3), we concatenate, i.e.
+$\parallel$, a sinusoidal positional encoding [37] $\mathbf{z}_{n}^{\mathbf{X}_{R}}\in\mathbb{R}^{256}$ to each patch embedding to encode spatial information.
+Next, we cluster the patches in $\hat{\mathbf{B}}^{\mathbf{X}_{R}}$ for each resolution $R$ using $k$-means clustering [24]:
 […]
-where $C_{k}$ is the set of $k$ clusters and $\mu_{i}$ are the cluster centroids.
-Running the clustering procedure with a 2 to 8 range for $k$ on two different image resolutions results in 14 different cluster assignments.
+where $C^{R}_{k}$ is the set of $k$ clusters for resolution $R$ and $\mu_{j}$ are the cluster centroids.
+Running the clustering procedure with $k\in\{2,\ldots,8\}$ on two different image resolutions results in 14 unique cluster assignments.
 Cross-clustering Consistency.
@@
 The probability $P(n|p)$ of $p$ being assigned to a particular cluster $n$ is defined as the relative frequency of $n$ among labels $\mathbf{L}(p)$:
-$P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }\mathbf{L}_{i}(p)=n\\ 0,&\text{\small else}\end{cases}$ (10)
+$P(n|p)=\frac{\sum_{i=1}^{m}\mathbbm{1}}{m},\text{\small with }\mathbbm{1}=\begin{cases}1,&\text{\small if }\mathbf{L}_{i}(p)=n\\ 0,&\text{\small else}\end{cases}.$ (10)
 The predictions of a single $k$-means predictor tend to contain high levels of noise, likely due to the high dimensionality of the embeddings.
@@
 Cluster Denoising.
-To further improve the locality and semantic meaningfulness of clustered feature representations, we apply a Conditional Random Field (CRF) [^2] We use the crfseg Python library.
-[13] and majority filter.
+To further improve the locality and semantic meaningfulness of clustered feature representations, we apply a Conditional Random Field (CRF) [14] and majority filter.
 CRF is a discriminative statistical method that is used to denoise predictions based on local interactions between them.
@@
 Our implementation is specifically tailored for refining 2D segmentation map, using a mean field approximation with a convolutional approach to iteratively adjust the probability distributions of each image patch’s cluster indices.
-Key to this process is the use of a Gaussian filter in the pairwise potentials, which ensures spatial smoothness and consistency in the segmentation.
+In the pairwise potentials, we use a Gaussian filter to ensure spatial smoothness and consistency in the segmentation.
 The application of the CRF yields embeddings which are less noisy and more cohesive than the original aligned $k$-means result.
@@
 Captioning.
-The next step involves turning clustered, denoised and enhanced BLIP embeddings into text.
+The next step involves turning clustered, denoised and enhanced embeddings into text.
 The BLIP text decoder is a transformer architecture capable of processing unordered sets of embeddings of arbitrary size.
-We leverage this feature and feed flattened subsets of patch embeddings, each corresponding to a cluster, to the BLIP text decoder.
+We leverage this feature and feed flattened subsets of patch embeddings, each corresponding to a cluster, to the text decoder.
 Spatial information is preserved due to the presence of positional embeddings added in the clustering step.
 With this technique, our method essentially infers semantic categories captured by clusters and represented by BLIP embeddings.
-To the best of our knowledge, we are the first to use the text decoder in this manner, enabling local captioning for which it was never specifically trained.
+To the best of our knowledge, we are the first to use the text decoder in this manner, enabling local captioning without specifically training for it.
 The caption generation is stochastic, with different object namings appearing in the captions depending on initialization.
-To obtain a rich, unbiased and diverse set of object names, we regenerate captions with each embedding multiple times in caption generation cycles.
-3.2 Caption Filtering
-The captions generated by the BLIP text decoder are sentences in natural language.
+To obtain a rich, unbiased and diverse set of object names, we regenerate captions with each embedding with multiple inference cycles.
+Caption Filtering.
+The captions generated in the previous step are sentences in natural language.
 For our task, we are only interested in the class names present in each sentence.
-To obtain these, we filter the sentence down to relevant class names using noun extraction.
-Captions are parsed using the spaCy model333www.spacy.io/models/en#en_core_web_sm to provide a part-of-speech label for each word in the caption.
-Words labeled as nouns are kept and converted into their singular form through lemmatization.
-We collect all nouns generated by different clusters and cycles into one noun set and remove any duplicates, as well as nouns which do not appear in the WordNet dictionary.
-3.3 Self-Guidance
-The output of the clustered BLIP embeddings (see Sect.
-3.1) is a 32x32 grid with cluster index assignments.
-Each element of the grid corresponds to an image patch from the original image, which enables us to create segmentation masks - defined as a union of areas covered by image patches with the same index - essentially for free.
-For example, Fig.
-3 shows the clustered output as a 2-dimensional mask, which succeeds in masking relevant objects to a certain extent.
-However, the segmentation outputs are constrained to a 32x32 resolution of the BLIP encoder.
-Upsampling yields unsatisfactory results, since objects appear oversegmented and boundaries remain unsharp.
-Another challenge is to associate each mask with one class name.
-In this cluster representation, masked areas are described with sentences rather than single nouns.
-When such a sentence contains multiple nouns, it is not trivial which noun to select as the predicted class name for that mask.
-To perform effective and accurate auto-vocabulary semantic segmentation, we leverage BCC’s strength in generating an elaborate set of relevant class names and use this as textual guidance for a pre-trained OVS model capable of producing high-resolution outputs.
-BCC is model-agnostic, therefore any kind of OVS model that expects an image and a set of class labels can be combined with our method.
-We focus on X-Decoder[42], which is a popular and well-performing OVS model.
-3.4 Custom Open-Vocabulary Evaluator
+To obtain these, we filter the sentence down to relevant class names by extracting all nouns using part-of-speech labels for each word in the caption.
+Nouns are kept and converted into their singular form through lemmatization.
+We collect all nouns generated by different clusters and cycles into one target vocabulary and remove any duplicates, as well as nouns which do not appear in the WordNet dictionary.
+3.2 Segmentation through Self-Guidance
+The output of the clustered BLIP embeddings is a 32x32 grid with cluster index assignments (see Sect.
+3.1), with each element corresponding to an image patch from the original image.
+This enables the extraction of segmentation masks - defined as a union of areas covered by image patches with the same index - essentially for free.
+For instance, Fig.3 and 4 show the clustered output as a 2D mask that partially captures relevant objects.
+However, the low resolution of the segmentations are unsatisfactory, and upsampling leads to oversegmented objects or unsharp boundaries.
+To perform effective and accurate auto-vocabulary semantic segmentation, we leverage BBoost’s strength in generating an elaborate set of relevant class names from clusters instead - and use this as textual guidance for a pre-trained OVS model capable of producing high-resolution outputs.
+BBoost is model-agnostic, allowing our approach to integrate with any OVS model that accepts an image and a set of class labels.
+In this work, we focus on X-Decoder [48], a popular and well-performing OVS model.
+4 Evaluation of Auto-Classes
 As discussed in Sect.
@@
 Hence, evaluation is possible by having access to the ground truth of those target class names during evaluation.
-However, in scenarios where the target class names are discovered rather than pre-specified, as in our framework, there may be a lack of direct alignment between the semantics of these categories and the classes used in the annotations.
-In zero-guidance segmentation [28], the authors have proposed to align class names based on the cosine-similarity in the latent space of Sentence-BERT [26] or CLIP [25].
-In our initial experiments, this yielded unsatisfactory results where, even in obvious cases, the open-vocabulary classes were linked to the incorrect class (e.g.
-taxi to road instead of car).
-This led to a decline in performance as measured by segmentation metrics, yet the method showed promising results from a qualitative perspective.
-Relations between two class names can be complex, such as synonymity, hyponymy, hypernymy, meronymy or holonymy, which exceed the capabilities of a cosine-similarity criterium in the respective latent spaces.
-As a possible solution, we propose an LLM-based Auto-Vocabulary Evaluator (LAVE).
-Specifically, we employ the Llama 2 7B Large Language model444www.huggingface.co/meta-llama/Llama-2-7b with the task of mapping predicted auto-vocabulary categories to target categories in the dataset.
-First, AutoSeg returns a segmentation mask with classes from an auto-vocabulary generated by BCC.
-Second, LAVE collects all uniquely predicted auto-vocabulary categories in the output and builds a mapper from each predicted category to the most relevant or similar target category.
-Thirdly, all pixel values in the segmentation mask belonging to each class are updated respectively according to the mapper.
-Finally, the mean Intersection over Union (mIoU) is computed with the updated segmentation mask.
-Pseudocode and prompts for LAVE are provided in the supplementary material.
+However, in scenarios where the target class names are discovered, as in our framework, rather than pre-specified there may be a lack of direct alignment between the semantics of these categories and the classes used in the annotations.
+In zero-guidance segmentation [32], the authors have proposed to align class names based on the cosine similarity in the latent space of Sentence-BERT [30] or CLIP [29].
+In our initial tests, however, this approach misaligned obvious class pairs (e.g., taxi mapped to road instead of car), thereby reducing segmentation accuracy despite promising qualitative results.
+Relations between two class names can be complex and ambigious [44], such as due to synonymity, hyponymy, or hypernymy, which exceed the capabilities of a cosine-similarity criterium in the latent space.
+To address this, we propose an LLM-based Auto-Vocabulary Evaluator (LAVE), leveraging the Llama-2-7B language model [36], to map predicted auto-vocabulary categories to target dataset classes.
+LAVE gathers all predicted auto-vocabulary classes and maps each category to the most relevant or similar class in the known vocabulary.
+It then updates all pixel values in the predicted segmentation masks according to the mapping, after which the mean Intersection over Union (mIoU) is calculated using the updated mask.
+Though not an integral part of our method, LAVE greatly reduces the mapping effort which is often infeasible to do manually, requiring $C\times A$ comparisons between $C$ known classes and $A$ auto-classes.
+Pseudocode and prompts for LAVE are included in the supplementary material.
 Figure 4: Segmentation with VLMs.
-Example outputs on PASCAL VOC/Context (top), ADE20K (middle) and Cityscapes (bottom) by feeding plain BLIP embeddings to X-Decoder, directly using BCC embeddings as masks and AutoSeg.
-Notably, our method segments images in the most comprehensive and accurate manner.
+Example outputs on PASCAL VOC/Context (top), ADE (middle) and Cityscapes (bottom) by (left to right) directly using BBoost embeddings as masks, feeding plain BLIP embeddings to X-Decoder or AutoSeg.
+Notably, our method segments images in the most comprehensive and semantically accurate manner.
+Table 1: Ablations on Segmentation with VLMs (c-mIoU). Leveraging BBoost semantic embeddings for segmentation with AutoSeg outperforms baselines on all four datasets.
+Method VOC [9] PC [27] ADE [47] CS [6]
+BBoost Embeddings 16.3 16.3 11.3 0.85
+X-Decoder [48] + BLIP [18] 38.0 35.3 26.7 29.2
+AutoSeg (Ours) 71.8 47.7 29.2 35.8
 5 Experiments
 5.1 Experimental Setup
-We evaluate our method on three popular semantic segmentation validation datasets: PASCAL VOC [9] and Context [23], ADE20K [41] and Cityscapes [6] with 20, 459, 847 and 20 classes respectively.
-These datasets cover a wide range of difficulty and class diversity.
-For example, ADE20K is challenging with its many and infrequently appearing classes, while Cityscapes contains many object instances per image.
-We list dataset statistics in Tab.
-2.
+Datasets.
+AutoSeg is evaluated on four popular semantic segmentation validation datasets: PASCAL VOC [9] and Context (PC) [27], ADE20K (ADE) [47] and Cityscapes (CS) [6] with 20, 459, 847 and 20 classes respectively, covering a wide range of difficulty and class diversity.
+For instance, ADE is challenging with its many and infrequently appearing classes, while CS contains many instances per image (see Tab.
+2 for detailed statistics).
 Evaluation.
@@
 As mentioned in Sec.
-3.4, class predictions in the segmentation mask are mapped using LAVE before the mIoU is computed.
+4, auto-class predictions in the segmentation mask are mapped using LAVE before the mIoU is computed.
 Furthermore, to study the class-agnostic segmentation performance of different VLMs, we report class-agnostic mean Intersection over Union (cmIoU).
@@
 The cmIoU is then determined by calculating the arithmetic mean of these highest IoU values across the dataset, which can be formulated as $\text{cmIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}_{\text{max}_{i}}$ where $N$ is the total number of ground-truth segments in the dataset.
-Table 1: Ablations on Segmentation with VLMs.
-Class-agnostic mIoU values for different variants of our method.
-AutoSeg, which combines BCC with X-Decoder segmentation, achieves the best performance on all four datasets.
-Approach VOC [9] PC-459 [23] ADE20K [41] Cityscapes [6]
-X-Decoder [42] + BLIP [16] 38.0 35.3 26.7 29.2
-Blip-Cluster-Caption 16.3 16.3 11.3 0.85
-AutoSeg 71.8 47.7 29.2 35.8
 Implementation details.
@@
 Parameters of the clustering include image scales encoded by BLIP ($384\times 384$ and $512\times 512$), the $k$ values of $k$-means clustering (2 to 8), the parameters of Gaussian smoothing in the CRF (smoothness weight of 6 and smoothness $\theta$ of 0.8), the number of iterations of majority filtering (8) and the feature dimension size of the positional embeddings (256).
-As for text generation, we opted for nucleus sampling with a minimum length of 4 tokens and maximum of 25, top P value of 1 and repetition penalty of 100 to ensure as many unique nouns as possible.
+For text generation, we use nucleus sampling with a minimum length of 4 tokens and maximum of 25, top P value of 1 and repetition penalty of 100 to ensure as many unique nouns as possible.
 These parameters were determined using Bayesian optimization on VOC with the goal of maximizing the mIoU.
+For CRF denoising and part-of-speech tagging, we use the crfseg [14] and spaCy [11] libraries respectively.
 5.2 Ablations
-To investigate the performance of different configurations of our framework, we present two ablation studies.
-The first ablation study investigates different methods of using a VLM to obtain segmentation masks.
-In the second, we examine the effect of different amounts of captioning cycles.
+To assess our framework we investigate various segmentation methods with captions, the impact of captioning cycles, generated/fixed vocabulary similarity, alternative mappers and the effects of cluster denoising.
+The latter three ablation studies are detailed in the supplementary material.
 Segmentation with Vision Language Models.
-A VLM can be leveraged in different ways to obtain segmentation masks.
-In this ablation study, we study the benefit of using an auto-vocabulary generated with BCC for segmentation as in AutoSeg.
-In order to do so, we compare our method against a scenario where plain BLIP embeddings are used for captioning instead of enhanced ones.
-Furthermore, we also evaluate the extent to which BCC embedding outputs themselves can be used directly as segmentation masks (see Sect.
-3.3) after upsampling.
-Due to the difficulty of establishing correspondence between BCC clusters and ground-truth labels, we evaluate the segmentation quality using cmIoU.
-Quantitative results are reported in Tab.
-1.
-On all four datasets, AutoSeg performs most effectively.
-Comparing X-Decoder + BLIP to AutoSeg qualitatively, one can observe that the former suffers from false positives and false negatives simultaneously, which in turn serves as an explanation for the performance gap in cmIoU.
-For example, in PASCAL, it incorrectly masks the area around the bus at the cost of segmenting people.
-In ADE20K and Cityscapes, it ignores large parts of the area which could appear to BLIP as insignificant.
-Our framework ensures that smaller, less salient objects are effectively captioned.
-As expected, intermediate masks of BCC do not segment objects sufficiently accurate due to the low resolution of the embedding space.
-Interestingly, it is still capable of segmenting global areas linked to certain objects in PASCAL and ADE20K, while failing completely for Cityscapes, likely due to the higher instance count and higher resolution of images.
-Table 2: Ablations on Captioning Cycles.
-The optimal number of captioning cycles varies per dataset.
-On Cityscapes, which has images in the highest resolution, most instances ($\overline{I}$) and classes on average ($\overline{C}$), AutoSeg benefits from 25 cycles, while between 1 and 15 captioning cycles for datasets with fewer instances per image suffices.
-                                 Number of Captioning Cycles
-Dataset $\overline{I}$ $\overline{C}$ $C$ 1 5 10 15 20 25
-PASCAL VOC [9] 3.5 3.5 20 85.3 84.4 87.1 85.7 83.3 83.3
-PASCAL Context [23] 18.9 6.2 459 10.1 9.7 11.2 11.7 9.6 10.8
-ADE20K [41] 19.5 10.5 847 5.9 5.8 5.1 5.2 5.1 5.6
-Cityscapes [6] 34.3 17 20 28.2 28.1 28.4 28.3 27.9 29.5
+Upsampled BBoost embeddings could potentially act as segmentation masks on their own given the clustering of semantic areas (see Sec.
+3.2).
+In this ablation study, we compare this baseline to AutoSeg which utilizes BBoost’s low-resolution semantics for high-resolution outputs.
+We assess segmentation quality using class-agnostic (c-mIoU), with results presented in Tab.
+1 and Fig.
+4.
+Across all datasets, AutoSeg performs best, with the baseline X-Decoder + BLIP showing issues of false positives and negatives that explain the performance gap.
+For example, it incorrectly segments areas around objects in VOC/PC (e.g.
+person for building) and ignores significant regions in ADE and CS.
+In contrast, AutoSeg effectively captures smaller objects and details.
+As expected, low-resolution BBoost masks are inaccurate but can still segment general areas in VOC/PC and ADE, though they struggle in CS due to higher object density.
+Table 2: Ablations on Captioning Cycles (mIoU) and Generated Classes.
+The optimal number of captioning cycles depends on the dataset, influenced by the image content and annotation method.
+Our method identifies a substantially larger variety of unique classes (Gen.) than those annotated by humans (Ann.), highlighting its capacity to capture finer context.
+$\overline{I}$ and $\overline{C}$ denote the average number of instances and classes per image.
+Number of Captioning Cycles Data Properties
+Dataset 1 5 10 15 20 25 Gen. Ann. $\overline{I}$ $\overline{C}$
+PASCAL VOC [9] 85.3 84.4 87.1 85.7 83.3 83.3 938 20 3.5 3.5
+PASCAL Context [27] 10.1 9.7 11.2 11.7 9.6 10.8 721 459 18.9 6.2
+ADE20K [47] 6.0 5.8 5.1 5.2 5.1 5.6 1578 847 19.5 10.5
+Cityscapes [6] 28.2 28.1 28.4 28.3 27.9 30.0 395 20 34.3 17
 Captioning Cycles.
-A key aspect of our model is its capability to iteratively repeat the captioning process within BCC, enabling it to generate a more exhaustive collection of nouns which accurately describe the scene.
-This approach proves especially beneficial in instances where objects are initially overlooked or described with less precise semantic terms.
-To assess the impact, we explore the effects of single versus multiple captioning iterations — 5, 10, 15, 20, or 25 times — using the mIoU metric.
+A key aspect of our model is its capability to repeatedly enrich the vocabulary with BBoost, enabling it to generate a more exhaustive collection of classes which accurately describe the scene.
+This approach proves especially beneficial for instances where objects are initially overlooked or described with less precise semantic terms.
+To assess its impact, we explore the effects of various captioning iterations using the mIoU metric.
 The findings, presented in Tab.
 2, reveal that the optimal number of captioning cycles varies significantly with the dataset in question.
-Specifically, we observe that for the Cityscapes dataset, characterized by its high average number of instances and unique classes per image (denoted as $\overline{I}$ and $\overline{C}$ in Tab.
+Specifically, we observe that for CS, characterized by its high number of instances and unique classes per image (denoted as $\overline{I}$ and $\overline{C}$ in Tab.
 2, respectively), a greater number of cycles yields the most benefit.
-For datasets with a lower average number of instances and unique classes per image, like VOC or ADE20K, fewer captioning cycles are required for BCC to identify all relevant class names.
-As anticipated, our model requires a greater number of captioning cycles for the Context dataset, which, despite featuring images identical to those in VOC, includes annotations for an additional 439 object categories.
-Remarkably, AutoSeg successfully handles ADE20K, the dataset with the largest number of unique classes in our experiments, using merely a single captioning cycle.
+For datasets with a lower average number of instances and unique classes per image, like VOC or ADE, fewer captioning cycles are required for BBoost to identify all relevant class names.
+As anticipated, our model requires a greater number of captioning cycles for the CS dataset given its additional 439 object categories.
+Remarkably, AutoSeg successfully handles ADE, the dataset with the largest number of unique classes in our experiments, using merely a single captioning cycle.
 This outcome underscores our model’s ability in efficiently addressing complex datasets which resemble an open-ended setting.
-Table 3: Quantitative Results (mIoU).
-AutoSeg performs competitive with existing OVS methods which require class names, while achieving superior performance on the zero label setting.
-Results with dashes are not published and unavailable.
-Method Auto-Classes VOC [9] PC-459 [23] ADE20K [41] Cityscapes [6]
+Finally, we note that through its captioning framework, AutoSeg perceives significantly more distinct classes compared to the hand-crafted fixed vocabularies.
+5.3 Quantitative Analysis
+Auto-Vocabulary Segmentation Setting.
+Our quantitative analysis begins by comparing AutoSeg with other auto-vocabulary methods, such as Zero-Guidance Segmentation (ZeroSeg [32]), as well as LLM-aided segmentation methods like LISA [16] with alternative captioning methods like LLaVA [21].
+We also evaluate a configuration combining BLIP and X-Decoder, where instance crops predicted by SAM [13] are used for captioning to simulate locality.
+We use the same caption filtering as in our method.
+The results, shown in Tab.
+3, indicate that our method outperforms ZeroSeg, the only other true auto-vocabulary approach.
+In contrast, LISA requires multiple inferences with individually prompted categories to achieve scene segmentation—a notably more complex and less efficient approach that also yields poor performance.
+Captioning based on smaller object crops is similarly ineffective compared to using enhanced vision-language features that maintain spatial locality.
+Moreover, while LISA provides LLM-guided captions, these lack the contextual specificity needed to establish an effective vocabulary for X-Decoder.
+Although AutoSeg also utilizes this segmentation backbone, it achieves a substantially higher performance.
+This comparison not only highlights the unique strengths of AutoSeg but also reveals the practical limitations of current LLM-aided approaches in segmenting scenes with multiple object categories.
+Table 3: Quantitative Auto-Vocabulary Results (mIoU).
+The wide availability of captioning and OVS methods allows various combinations to design an auto-vocabulary method, but they are not necessarily performing well.
+AutoSeg performs superior over the only Auto-Vocabulary method ZeroSeg, as well as over AVS-adapted methods with alternative captioners or segmentors.
+Auto-Vocabulary Segmentation Method VOC [9] PC [27] CS [6]
+LLaVA + LISA [21, 16] 7.7 0.2 1.5
+ZeroSeg [32] 20.1 11.4 -
+SAM + BLIP + X-Decoder [13, 18, 48] 41.1 11.3 27.4
+LLaVA + X-Decoder [21, 48] 56.7 11.4 23.4
+AutoSeg (Ours) 87.1 11.7 30.0
+Figure 5: Qualitative Results.
+AutoSeg shows remarkable capability to identify out-of-vocabulary categories, such as hawk or coke, and segment them accurately across different datasets.
+Images are from the VOC/PC, ADE and CS datasets.
+Table 4: Open/Auto-Vocabulary State of the Art Comparison (mIoU).
+AutoSeg surpasses ZeroSeg [32] in human-free segmentation and remains competitive with some well-known OVS methods that rely on human input, while outperforming others (3 out of 8 methods on VOC and 3 out of 7 methods on PC).
+Results with dashes indicate unpublished or unavailable data.
+Unknown VOC [9] PC [27] ADE [47] CS [6]
+Method Vocabulary (20) (459) (847) (20)
+OVS Segmentation methods with prompted (known = ground truth) vocabulary
 LSeg [17] 47.4 - - -
-OVSegmentor [36] 53.8 20.4 - -
-SimSeg [39] 57.4 26.2 - -
-OVDiff [11] 69.0 31.4 - -
-OpenSeg [10] 72.2 45.9 8.8 -
-OVSeg [17] 94.5 55.7 9.0 -
-ODISE [34] 82.7 57.3 11.0 -
-SAN [37] 94.6 57.7 12.4 -
-CAT-Seg [5] 97.2 62.0 13.3 -
-ZSSeg [38] 83.5 47.7 7.0 34.5
-X-Decoder [42] 96.2 65.1 6.4 50.8
-FC-CLIP [40] 95.4 58.4 14.8 56.2
+OpenSeg [10] 72.2 9.0 8.8 -
+OVSeg [19] 94.5 11.0 9.0 -
+ODISE [41] 82.7 13.8 11.0 -
+SAN [43] 94.6 12.6 12.4 -
+CAT-Seg [5] 97.2 19.0 13.3 -
+X-Decoder [48] 96.2 16.1 6.4 50.8
+FC-CLIP [46] 95.4 12.8 14.8 56.2
+AVS Segmentation methods with auto-generated (unknown) vocabulary
 ZeroSeg [32] 20.1 11.4 - -
-AutoSeg (Ours) 87.1 11.7 5.9 29.5
-4.3 Quantitative Analysis
-We compare our method against existing OVS methods that require class names to be given, as well as with Zero-Guidance Segmentation (referred to as ZeroSeg [28]).
+AutoSeg + LAVE mapper (Ours) 87.1 11.7 6.0 30.0
+AutoSeg + Manual mapper (Ours) 88.2 12.8 6.2 31.1
+Open-Vocabulary Segmentation Setting.
+We compare our method against existing OVS methods that require class names to be given.
 Tab.
 4 shows the results.
-Without any specification of class names through user input, AutoSeg achieves 86%, 18%, 40% and 52% of the best performing OVS methods on VOC, Context, ADE20K and Cityscapes respectively.
-The difference in performance can be attributed to AutoSeg needing to identify the subjects for segmentation independently, while OVS methods are explicitly instructed on the specific objects to segment.
-Remarkably, however, AutoSeg is able to remarkably surpass the performance many OVS methods on VOC, such as OVDiff [11], OpenSeg [10] or ODISE [34], without any access to the ground truth class names.
-While still leaving room for improvement, our model compares competitively with OVS methods on ADE20K and Cityscapes, two very challenging datasets either high in number of unique classes or instances.
-Comparing to ZeroSeg, AutoSeg achieves superior performance on VOC (87.1 over 20.1 mIoU) and Context (11.7 over 11.4 mIoU).
-Furthermore, we set a new state-of-the-art on ADE20K and Cityscapes under the auto-class generation task setting.
-This outcome underscores the efficacy of our method in dealing with complex scenes that have an open-ended nature, such as ADE20K with its large number of rare classes.
-While other models have to decide between 847 possibilities, AutoSeg suggests a smaller yet more relevant set of class names.
+In addition to the results mapped with LAVE, we provide results with one manual mapper per dataset (note that we provide a manual mapper for this table only, given the significant manual effort to construct it).
+Without any specification of class names through user input, AutoSeg matches 91%, 67%, 42% and 55% of the best OVS method performance on VOC, PC, ADE and CS respectively.
+It should be noted that this metric reflects the performance on the known, annotated classes, while additional open-ended classes are potentially mapped.
+Despite not being explicitly instructed with the known vocabulary contrary to OVS methods, AutoSeg is remarkably able to surpass six of them on VOC and PC, such as OpenSeg [10], ODISE [41] or OVSeg [19].
+While still leaving room for improvement, our model compares competitively with OVS methods on ADE and CS, two very challenging datasets either high in number of unique classes or instances.
+Compared to ZeroSeg, AutoSeg achieves superior performance on VOC (88.2 over 20.1 mIoU) and PC (12.8 over 11.4 mIoU).
+Furthermore, we set the first benchmark on ADE and CS under the unknown vocabulary task setting.
+This outcome underscores the efficacy of our method in dealing with complex scenes that are open-ended in nature, such as ADE with its large number of rare classes.
+Finally, results obtained with LAVE mappings show little difference with manually constructed ones.
+This indicates that at marginal cost, LAVE can act as a feasible bridge between a known and unknown vocabulary.
 Our method deals well with high numbers of instances and can successfully identify various non-salient objects.
-A comparison with the findings from Tab.
-2 reveals that while BCC struggles with providing high-quality masks on its own, it is able to identify the accurate target classes, which can subsequently be segmented with precision.
-Additional experiments are provided in the supplementary material, in which we generate class names with other image captioning methods and compare the segmentation performance of such vocabularies.
-Overall, our results demonstrate that the integration of BCC and OVS harnesses the strengths of both, leading to enhanced open-ended recognition capabilities.
-Figure 5: Qualitative Results.
-AutoSeg shows remarkable capability to identify out-of-vocabulary categories, such as hawk or coke, and segment them accurately across different datasets.
-The first two rows highlight results on from PASCAL VOC/Context, while the bottom two rows are from ADE20K and Cityscapes respectively.
+While BBoost struggles with producing high-quality masks by itself (as seen in Tab.
+2), it is highly effective in embedding target classes accurately which can be segmented with precision afterwards.
+Overall, our results demonstrate that the integration of BBoost with OVS harnesses the strengths of both, leading to enhanced open-ended recognition capabilities.
 5.4 Qualitative Analysis
 Fig.
-5 displays qualitative results on the four datasets, where AutoSeg demonstrates its capability to comprehensively detect and accurately segment relevant object categories within the images.
-Remarkably, it predicts segmentation masks for specific class names such as hawk, piano, wildflower, officer or coke - all categories not listed in the annotations of their respective datasets.
-Moreover, in certain instances, AutoSeg successfully captures additional contextual details, such as graduation or exercise, illustrating the model’s capabilities at recognizing objects in scenes in a genuinely open-ended manner.
-Additional results, as well as failure cases, are available in the supplementary material.
+5 displays qualitative results on the four datasets, where AutoSeg demonstrates its capability to accurately detect and segment relevant object categories present in the images but missing from the ground truth vocabulary.
+Remarkably, it predicts segmentation masks for classes such as moped, presentation, coke, courtroom or hawk.
+Moreover, in certain instances, AutoSeg successfully captures additional contextual details, such as graduation or reflection, illustrating the model’s capabilities at semantic segmentation in a genuinely open-ended manner.
+Additional results, including failure cases, are included in the supplementary material.
 6 Conclusion
-In this work we introduced AutoSeg, a novel method which leverages a vision-language model to automatically identify relevant target classes and segment them - a task we termed Auto-Vocabulary Semantic Segmentation.
-Additionally, we proposed LAVE, a new evaluation method which employs an LLM to map open-ended class names to ground-truth labels.
-AutoSeg demonstrates remarkable open-ended recognition capabilities, achieving state-of-the-art performance on the zero label setting, while being competitive with open-vocabulary segmentation models which require ground-truth label specification.
+This paper introduced AutoSeg, a novel method which leverages a vision-language model to automatically generate relevant target classes and segment them.
+Additionally, we proposed LAVE, a new evaluation framework which maps open-ended class names to ground-truth labels.
+AutoSeg shows open-ended recognition capabilities, achieving state-of-the-art performance in the zero label setting, while being competitive with open-vocabulary segmentation models which require provided ground-truth labels.
 Acknowledgements.
 This work was supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.
-Supplementary Material
-Appendix A. LLM-based Auto-Vocabulary Evaluator
-This appendix provides more insight into the pseudocode and text prompts of the LLM Mapper which maps auto-vocabulary classes to the fixed-vocabulary classes present in the annotations of each respective dataset.
-The LLM Mapper is a key component of our LLM-based Auto-Vocabulary Evaluation Metric (LAVE) and enables the evaluation of predicted auto-vocabulary classes on public benchmarks.
-By doings so, we enable comparison with with open-vocabulary methods, such as in Tab.
-3, which predict a fixed-vocabulary for evaluation.
-LLM Mapper Pseudocode
-The LLM Mapper is tasked with creating a mapping dictionary which maps from auto-vocabulary classes to fixed-vocabulary classes.
-The resulting mapping dictionary can be used during inference to map pixel-level class predictions belonging to an automatically generated - out-of-vocabulary- class to fixed-vocabulary class indices.
-In Alg.
-1, we detail the procedure to create the mapping dictionary.
-LLM Mapping Prompts
-The generateDialogs() function in the LLM Mapper takes a prompt template as one of its inputs.
-This prompt template is dependent on the dataset and is used to generate dialogs to query the LLM with.
-For PASCAL VOC and Cityscapes, we explicitly specify the list of categories given its short length with 20 classes each.
-For PASCAL Context and ADE20K, with its 459 and 849 classes respectively, we rely on the LLM’s knowledge of the datasets.
-Explicitly naming each object class from these datasets causes the LLM to run into memory issues when queried with long dialogues.
-We specify the prompt templates for each respective dataset as follows:
-PASCAL VOC and Cityscapes
-To which class in the list <dataset> is <name> exclusively most
-similar to? If <name> is not similar to any class in the list or
-if the term describes stuff instead of things, answer with
-‘background’. Reply in single quotation marks with the class name
-that is part of the list <dataset> and do not link it to any other
-class name which is not part of the given list or ‘background’.
-where <noun> is the noun to map to the vocabulary and <dataset> is the list of classes of PASCAL VOC or Cityscapes.
-PASCAL Context and ADE20K
-To which class in the <dataset> dataset is <name> exclusively most
-similar to? If <name> is not similar to any class in <dataset> or
-if the term describes stuff instead of things, answer with
-‘background’. Reply in single quotation marks with the class name
-that is part of <dataset> and do not link it to any other class
-name which is not part of the given list or ‘background’.
-where <noun> is the noun to map to the vocabulary and <dataset> is either ‘PASCAL-Context 459’ or ‘ADE20K’.
-Appendix B. Alternative Class Generation Baselines
-In Tab.
-4, we provide additional comparisons with alternative methods to generate class names from images.
-In particular, we compare against an altered version of BLIP to caption the image at an object instance level with segments retrieved from the Segment Anything Model (SAM[12].
-We use segment everything prompting, i.e.
-the model’s object proposal generation functionality.
-Furthermore, we compare against Large Language and Vision Assistant (LLaVa[18], generating classes with the prompt ‘Please name all the elements present in the image as a list of nouns’.
-We compare the mIoU performance on VOC, Context and Cityscapes and find that AutoSeg performs favorably over other approaches that are adjusted to segment with automatically generated classes.
-The performance gap is especially high for datasets with a small vocabulary, such as VOC and Cityscapes, possibly indicating that AutoSeg is better able in generating classes that are relevant for prediction while the other methods generate redundant classes which become false positive predictions.
-In contrast to both SAM and LLaVa, AutoSeg has an additional benefit of being independent of prompting.
-For this reason, our method does not require prompt engineering in order to obtain class names effective for accurate segmentation.
-Table 4: Experiments with Alternative Class Generation Baselines (mIoU).
-Experiments with alternative methods to generate classes showcase that AutoSeg is favorable for AVS across all three datasets, with larger performance gaps on VOC and Cityscapes with a smaller vocabulary size.
-Approach VOC [9] PC-459 [23] Cityscapes [6]
-SAM + BLIP + X-Decoder 41.1 11.3 27.4
-LLAVA + X-Decoder 56.7 11.4 23.4
-AutoSeg 87.1 11.7 29.5
-Appendix C. Visualization of Embeddings in BCC
-In Fig.
-6, we visualize intermediate 32x32 outputs of BCC to better understand what each individual step of Clustering, Alignment or Denoising achieves.
-Firstly, Fig.
-6 reveals that clustering BLIP embeddings results grouping pixels into semantic areas.
-With $k=2$, the image is divided an area which includes the cows (in blue) and another for everything else (green), while for $k=8$, various other objects - such as the pole in purple - get included.
-At this stage, the embeddings appear noisy and inconsistent across all $k$, which is why the combination of all clusters is so important.
-The alignment output shows that certain objects from a higher $k$ embeddings are now included in those of lower $k$ embeddings, i.e.
-the aligned embedding captures a more extensive list of objects present in the image.
-Finally, to further reduce the noise in the embedding, we apply CRF and a majority filter.
-We see that the final embedding consists of three unique semantic areas, which are eventually fed to the BLIP text decoder for captioning.
-Figure 6: Visualization of Intermediate BCC Outputs. Intermediate BCC outputs show how BLIP embeddings are enhanced with each individual step of Clustering, Alignment or Denoising.
-Appendix D. Additional Qualitative Results
-In Fig. 7,  8 and  9, we provide additional qualitative results of AutoSeg and discuss observed capabilities or limitations in the caption of each figure.
-(a) Success Case. AutoSeg can infer detailed class names, such as courtroom, which is non-existent in the extensive vocabulary of the dataset.
-(b) Success Case. The class badminton is not part of ADE20K. Despite taking up a very small area in the image, AutoSeg is still capable of segmenting it.
-(c) Success Case.
-AutoSeg displays advanced scene understanding capabilities.
-Here, it recognizes the out-of-vocabulary class reflection, which requires the model to reason about other auto-generated classes and structural properties in the image.
-(d) Success Case. The rare and out-of-vocabulary class name dinosaur is generated automatically and segmented by AutoSeg.. This is another example of the method’s efficacy on long-tailed classes.
-(e) Failure Case. AutoSeg’s occasionally struggles with indoor scenes from specific buildings, such as a mosque in this image. Our model mistakenly regards it a temple.
-(f) Failure Case.
-Here, AutoSeg predicts the zebra rug as a zebra as a result of noun parsing.
-After BCC has generated captions from the image, both zebra and rug get parsed into individual nouns.
-When fed to the segmentor, zebra eventually ends up being predicted as the more relevant class name.
-Figure 7: Additional Qualitative Results on ADE20K.
-(a) Success Case.
-Despite the visual similarities between a moped and motorcycle, AutoSeg correctly considers both and segments them accurately.
-A supervised model would fail to generalize to moped, as it is out of Cityscapes’ vocabulary.
-(b) Success Case. While annotations in Cityscapes consider all car types to be equal, AutoSeg often generates more specific descriptions, such as the SUV in this image.
-(c) Success Case.
-AutoSeg is able to provide class names which can be crucial for real-life scenarios, such as hazard in this example.
-The standard 20 classes in Cityscapes do not account for such objects descriptions.
-(d) Failure Case.
-Occasionally, generated class names, such as street, dominate the search space of the segmentor and cause it to attend more on that class.
-This may result in oversegmentation and other class names, such as car, to be ignored.
-Figure 8: Additional Qualitative Results on Cityscapes.
-(a) Success Case. AutoSeg is able to accurately classify and segment classes not present in the dataset, such as the ox.
-(b) Failure Case.
-Generated nouns are occasionally hyponyms or hypernyms of each other, causing the classes to “compete" for the final segmentation.
-Future work could possibly address this by checking for such relations between nouns and correcting for them.
-Figure 9: Additional Qualitative Results on PASCAL.
-Algorithm 1 LLM Mapper
-0:  nouns, dataset_vocabulary, llm_batch_size, prompt_template, LLM
-  all_responses, map_dict, skipped $\leftarrow$ initialize()
-  for batch in split(nouns, llm_batch_size) do
-     dialogs $\leftarrow$ generateDialogs(batch, prompt_template)
-     batch_responses $\leftarrow$ LLM(dialogs)
-     store(all_responses, (batch_responses, batch))
-  end for
-  for (batch_response, batch) in all_responses do
-     for (response, noun) in (batch_response, batch) do
-        answer $\leftarrow$ parseResponse(response)
-        common $\leftarrow$ intersect(answer, dataset_vocabulary)
-        updateDict(map_dict, noun, common)
-        updateSkipped(skipped, noun, common)
-     end for
-  end for
-  for key in map_dict do
-     if key is in vocabulary then
-        map_dict[key] $\leftarrow$ key
-     end if
-     if key is in skipped then
-        skipped $\leftarrow$ removeItem(skipped, key)
-     end if
-  end for
-  for skipped_key in skipped do
-     map_dict[skipped_key] $\leftarrow$ “background"
-  end for
-  return  map_dict